### v2

# Topological Phases of Many-Body Localized Systems: Beyond Eigenstate Order

Abstract
Many-body localization (MBL) lends remarkable robustness to nonequilibrium phases of matter. Such phases can show topological and symmetry breaking order in their ground and excited states, but they may also belong to an anomalous localized topological phase (ALT phase). All eigenstates in an ALT phase are trivial, in that they can be deformed to product states, but the entire Hamiltonian cannot be deformed to a trivial localized model without going through a delocalization transition. Using a correspondence between MBL phases with short-ranged entanglement and locality preserving unitaries—called quantum cellular automata (QCA)—we reduce the classification of ALT phases to that of QCA. This method extends to periodically (Floquet) and quasiperiodically driven ALT phases, and captures anomalous Floquet phases within the same framework as static phases. The QCA framework further generalizes to include symmetry-enriched ALT phases (SALT phases), and provides a large class of soluble models suitable for realization in quantum simulators. In systematizing the study of ALT phases, we both greatly extend the classification of interacting nonequilibrium systems and clarify a confusion in the literature which implicitly equates nontrivial Hamiltonians with nontrivial ground states.
Many-body localization (MBL) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] can protect robust, topological phases even in systems which are far from equilibrium [12, 13]. For instance, localized systems can support spontaneous symmetry breaking [12, 14, 15, 13], topological order [16, 17, 18, 19, 20, 21], and nontrivial phases with external driving [22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32], including time crystals [33, 34, 35, 36, 37, 38, 39, 40]. In fact, any Hamiltonian built out of commuting terms, of the kind frequently employed to study topological order [41, 42, 43, 44, 45, 46, 47, 48], becomes robustly localized by choosing strongly disordered coefficients for the commuting terms [20].
Two MBL systems that can be deformed into one another without going through a delocalization transition belong to the same localized phase [24, 23, 49, 25, 29, 50, 31]. Two Hamiltonians certainly belong to distinct localized phases when they support distinct eigenstate order [12, 14, 15, 13]—their eigenstates cannot be deformed into one another without their correlation length diverging. However, in this Letter, we argue eigenstate order is not the right concept to obtain a general understanding of localized topological phases. For one thing, not all ground state topological phases are compatible with MBL. More importantly, there exist anomalous localized topological phases [^1] The label “anomalous” descends from Refs. [22, 24]. It is not related to other usages of the word “anomalous” in the context of topological phases of matter. (ALT phases) [22, 24, 23, 49, 31, 50]. ALT phases have no eigenstate order, but nonetheless cannot be deformed to a trivial model—in which the commuting terms have disjoint support [51]—without going through a delocalization transition.
The existence of ALT phases in driven systems is recognized [22, 52, 23, 49, 31, 53, 27, 26, 54], but in static systems, ALT phases have only been discussed in the single-particle context [50]. Nonetheless, we propose that the static interacting models of Refs. [47, 48, 55, 56, 57] in fact represent ALT phases upon disordering the Hamiltonian.
Figure 1: (a) Many-body localization is characterized by commuting LIOMs, $\tau^{z}_{j}$ [6]. The LIOMs of an MBL model with short-ranged entanglement can be prepared from trivial, single-site LIOMs $\sigma^{z}_{j}$ by a QCA, $\tau^{z}_{j}=V\sigma^{z}_{j}V^{\dagger}$. ALT phases cannot be deformed back to a trivial model without delocalizing, even though they have no eigenstate order. (b) Multi-tone driven ALT phases correspond to loops or tori of QCA. The $\Omega$-spectrum property $\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}$ (5) implies that $n$-tone-driven ALT phases in $d$ dimensions correspond to $(d-1)$-dimensional phases with $n-1$ drives. The translation QCA in $d=1$ [58] corresponds with the anomalous Floquet phase in $d=2$ [23], which in turn corresponds with a two-tone-driven phase in $d=3$ with circulating surface states.
This Letter systematizes the study of ALT phases in strongly interacting systems. Specifically, we focus on localized topological phases satisfying a condition we refer to as short-ranged entanglement. We will argue that instead of eigenstate order, the right way to characterize all topological localized phases with short-ranged entanglement is via quantum cellular automata (QCA) [58, 59, 60, 56]—the locality preserving unitaries which map physical degrees of freedom to local integrals of motion (LIOMs) [Fig. 1(a)].
As QCA in spin systems have been classified in one [58] and two [61] dimensions, this correspondence allows for a classification of ALT phases in low dimensional systems. The same formalism extends to periodically (Floquet [62]) and quasiperiodically driven ALT phases [31], which are classified by loops or tori of QCA, respectively. Thus, we can treat static and $n$-tone-driven ALT phases in $d$ dimensions on the same footing. We also considerably develop the study of the topology of QCA, which allows us to classify all ALT phases with $d-n\leq 2$ [Fig. 1(b)].
Symmetry-enriched ALT phases (SALT phases) can also be classified by QCA, and we compute the classification for any symmetry group $G$ and $d-n\leq 1$. Our methods naturally provide explicit soluble models of driven (S)ALT phases which, being finite depth circuits, are well suited to realization in quantum simulators [63, 64, 65].
ALT phases.—We aim to classify many-body localized phases. A Hamiltonian $H$ is many-body localized when it possesses a complete set of commuting LIOMs $\tau^{z}_{j}$ such that [6]
$H=\sum_{j}h_{j}\tau^{z}_{j}+\sum_{jk}h_{jk}\tau^{z}_{j}\tau^{z}_{k}+\cdots.$ (1)
Here, $j$ indexes LIOMs, $\cdots$” indicates higher-body interactions, $h_{jk\ldots l}$ are real numbers which decay with the maximum distance between any of $\tau^{z}_{j},\ldots,\tau^{z}_{l}$, and we consider local degrees of freedom on the integer lattice in $d$ dimensions, ${\mathbb{Z}}^{d}$.
Equation (1) identifies MBL Hamiltonians with commuting models, which are independently used as exactly solvable models of topological order [41, 42, 43, 44, 45] and symmetry-protected topological (SPT) phases [16, 66, 67, 68, 46, 69, 48, 70]. In the MBL context, the LIOMs may have exponential tails.
MBL arising from strong random disorder is believed to be asymptotically unstable for $d>1$ [71], and doubt has arisen regarding the LIOM description even for $d=1$ [72, 73, 74, 75, 76, 77, 11]. The case of correlated disorder remains open [78, 79, 80, 81, 82]. Regardless, the Hamiltonian (1) continues to function as an effective model for astronomical prethermal timescales [72, 76, 77, 83, 10, 11]. Indeed, the timescale for topological response in a localized model tends to be parametrically separate from the thermalization time [31, 50, 23, 32, 25, 29, 30]. This justifies treating the topological classification of LIOM Hamiltonians (1) and the eventual thermalization process as separate problems. This Letter addresses the former.
We say two MBL Hamiltonians $H_{0}$ and $H_{1}$ are in the same localized topological phase if there is a continuous path of MBL Hamiltonians $H_{t}$ ($t\in[0,1]$) which connects them. Ignoring long-range resonances in the spectrum of $H_{t}$, it is possible to construct a continuous path of locality preserving unitaries $V_{t}$ (that is, $V_{t}AV_{t}^{\dagger}$ is localized near the support of $A$ for all operators $A$) such that $V_{t}$ conjugates the LIOMs of $H_{0}$ into LIOMs of $H_{t}$ [84, 85, 86, 87, 88, 89]. Indeed, this unitary simply follows the adiabatic evolution of the many-body eigenstates. Including long-range resonances prevents an adiabatic construction of $V_{t}$ [90], but it is expected that sufficiently dilute resonances can be incorporated such that there still exists a $V_{1}$ that conjugates LIOMs of $H_{0}$ into those of $H_{1}$ and is continuously connected to the identity [9]. Thus, if $H_{0}$ and $H_{1}$ are in the same localized phase, their LIOMs are related by a continuous path of locality-preserving unitary operators. Locality-preserving unitaries are also called quantum cellular automata (QCA), and we will use this terminology in the rest of the paper.
An immediate consequence is that if $H_{0}$ and $H_{1}$ have distinct eigenstate order, they cannot be in the same localized phase. However, the converse is not true: even if they have the same eigenstate order, it need not be possible to deform $H_{0}$ to $H_{1}$. This phenomenon is exemplified by ALT phases, which have no eigenstate order, but which cannot be deformed to a trivial MBL model with LIOMs supported on a single lattice site. We denote these trivial LIOMs $\sigma^{z}_{j}$, even when the local Hilbert space dimension is not 2.
An example of a nontrivial ALT phase is the three-dimensional 3-fermion Walker-Wang model [47, 48, 56]. In the absence of symmetry, this model is believed to have no eigenstate order [48, 91], but nonetheless it supports topologically ordered surface states protected by localization. Phrased in the language of MBL, [56, Theorem II.4] showed that the 3-fermion model, and, in fact, any MBL Hamiltonian $H$ with short-ranged entanglement—there are local unitaries $\tau^{x}_{j}$ which create and destroy excitations of $\tau^{z}_{j}$ without affecting other $\tau^{z}_{k}$ (Appendix A.1)—can be prepared from a trivial model by a QCA. That is, there is a QCA $V$ such that $H=VH_{\mathrm{trivial}}V^{\dagger}$ and $H_{\mathrm{trivial}}$ has single-site LIOMs $\sigma^{z}_{j}$ (Appendix A.2). We note that the property we call short-range entanglement for MBL Hamiltonians implies all eigenstates are short-range entangled in the usual sense.
Hence, short-range entangled MBL models can be parameterized by QCA as $VH_{\mathrm{trivial}}V^{\dagger}$. Paths of nondegenerate MBL Hamiltonians induce paths of unitaries as described above, so Hamiltonians in the same phase can be prepared by QCA which are connected by a path. Thus, the path-connected components of QCA can be used to classify phases of MBL Hamiltonians. An additional subtlety is that this parameterization is not unique—any QCA $V$ and $VP$ which differ by a permutation of the sites $P$ prepare identical LIOMs, and so represent Hamiltonians in the same phase (Appendix A.3).
We further allow stabilization by the stacking (tensor product) of additional ancillae—$V_{0}$ and $V_{0}\otimes\mathbbm{1}$ are said to be stably equivalent, and are considered to prepare Hamiltonians in the same phase [58, 51, 92, 93]. The space of stable equivalence classes of $d$-dimensional QCA is denoted $\mathrm{QCA}_{d}$, and its connected components $\pi_{0}(\mathrm{QCA}_{d})$ form an abelian group under tensor products of QCA [58, 94]. Short-range entangled localized phases are classified by the quotient of this group by permutation QCA.
The QCA formalism also lets us straightforwardly identify which phases are anomalous. For any simultaneous eigenstate $|0\rangle$ of all $\sigma^{z}_{j}$, $V|0\rangle$ is an eigenstate of $H$. This defines a map $e_{0}(V)=V|0\rangle$ from $\mathrm{QCA}_{d}$ to the space of invertible states [95] (short-range entangled states) in $d$-dimensions, $\mathrm{inv}_{d}$. There is an induced homomorphism
$e^{*}_{0}:\pi_{0}(\mathrm{QCA}_{d})\to\pi_{0}(\mathrm{inv}_{d}),$ (2)
where addition of states is again defined by stacking. The crucial observation is that the homomorphism $e_{0}^{*}$ is, in general, neither surjective nor injective. The lack of surjectivity means that not all invertible states can arise as eigenstates of commuting models (for instance, chiral states [96, 97]). The lack of injectivity (even accounting for equivalence by permutations) precisely means that nontrivial ALT phases exist. Indeed, ALT phases may be identified as the nontrivial phases in the kernel of $e^{*}_{0}$.
Most examples of ALT phases in the literature are driven systems [24, 23, 49, 25, 29, 50, 31]. We specialize to the case of multi-tone time dependence [98, 28, 99] (which includes Floquet driving as a special case), such that the Hamiltonian can be parameterized $H(t)=H(\theta_{1}(t),\ldots,\theta_{n}(t))$, where $\theta_{k}(t)=\omega_{k}t$ is a drive phase defined modulo $2\pi$. It is convenient to assemble the drive phases into a vector ${\vec{\theta}}_{t}=(\theta_{1}(t),\ldots,\theta_{n}(t))$. Then we say that the system is MBL with short-ranged entanglement if it has a locality preserving generalized Floquet decomposition ($\hbar=1$[28, 99]
$U(t)=V({\vec{\theta}}_{t})\exp\bigg{[}-it\bigg{(}\sum_{j}h_{j}\sigma^{z}_{j}+\cdots\bigg{)}\bigg{]}V({\vec{\theta}}_{0})^{\dagger}.$ (3)
Here, $U(t)$ is the time evolution operator. Thus we have a correspondence between multi-tone driven localized phases and parameterized families of QCA, $V({\vec{\theta}})$. That is, homotopy classes of continuous maps from the $n$-dimensional torus, ${\mathbb{T}}^{n}:=[{\mathbb{R}}/(2\pi{\mathbb{Z}})]^{n}$, to $\mathrm{QCA}_{d}$, which we denote $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$. Ignoring weak invariants [100, 101] corresponding to low dimensional cross sections of ${\mathbb{T}}^{n}$, we may instead consider the $n$th homotopy group $\pi_{n}(\mathrm{QCA}_{d})$.
Once more, there is an induced homomorphism $e^{*}_{0}:\pi_{n}(\mathrm{QCA}_{d})\to\pi_{n}(\mathrm{inv}_{d})$, and ALT phases are identified as the nontrivial elements of $\ker e^{*}_{0}$. As in the case of static systems, the parameterization by QCA is not unique, so that the classification of ALT phases is properly a quotient of $\ker e^{*}_{0}$. Further details regarding driven MBL phases are provided throughout Appendix A.
Classification.—The classification of static ALT phases in $d$ dimensions is determined by the classification of QCA. That is, by $\pi_{0}(\mathrm{QCA}_{d})$. There has been great recent progress in classifying QCA in spin systems—a complete classification now exists in one [58] and two [61] dimensions, and several nontrivial phases are known in three dimensions [56, 55, 57]. The classification of ALT phases inherits this progress.
For Floquet systems, there is another correspondence known between localized phases and QCA. Reference [23] showed that a periodically driven localized phase can be classified by its action on the edge when placed on open boundary conditions. This action is described by a QCA of one dimension lower, $\pi_{0}(\mathrm{QCA}_{d-1})$. On the other hand, both our construction and an equivalent formulation of unitary loops [49] show that Floquet localized phases should be classified by $\pi_{1}(\mathrm{QCA}_{d})$. We deduce that these groups must be isomorphic,
$\pi_{0}(\mathrm{QCA}_{d-1})\cong\pi_{1}(\mathrm{QCA}_{d}).$ (4)
The relationship Eq. (4) is reminiscent of a widely held conjecture that invertible states form an $\Omega$-spectrum [96, 95, 102, 103, 104]. That is, there is a homotopy equivalence [105] $\mathrm{inv}_{d-1}\simeq\Omega\mathrm{inv}_{d}$, where $\Omega X$ is the space of loops that begin and end at a specified base point in the space $X$. In particular $\pi_{n-1}(\mathrm{inv}_{d-1})\cong\pi_{n}(\mathrm{inv}_{d})$ for all $n$. We make the conjecture that QCA also forms an $\Omega$-spectrum,
$\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}\implies\pi_{n-1}(\mathrm{QCA}_{d-1})\cong\pi_{n}(\mathrm{QCA}_{d}).$ (5)
The base point for the loops in $\Omega\mathrm{QCA}_{d}$ is the identity.
Figure 2: The swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ maps a QCA $v$ to a loop of QCA $S(v,t)$ in one higher dimension. $S(v,t)$ is a finite depth circuit built from $v$, $v^{\dagger}$ (squares), and a path from the identity to a swap gate, $\mathrm{SWAP}(t)$ (line with crosses).
Equation (5) is interpreted through a picture of pumping [95, 104, 103]. Taking a loop $\gamma\in\Omega\mathrm{QCA}_{d}$ and restricting it to open boundary conditions, the action in the bulk must still be trivial at the end of the loop, due to the finite range of the QCA. However, on the $(d-1)$-dimensional edge, there may be a nontrivial action by some $v\in\mathrm{QCA}_{d-1}$ [23]. We say that $\gamma$ pumps $v$ across the system, and Eq. (5) hypothesizes that the homotopy structure of the loop space $\Omega\mathrm{QCA}_{d}$ is completely characterized by the pumped QCA $v$.
More formally, we argue that the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is a homotopy equivalence. Given some $v\in\mathrm{QCA}_{d-1}$, the swindle $S(v,t)$ is defined through a finite depth quantum circuit [95, 106, 60, 59] (Fig. 2). Let $\mathrm{swap}_{jk}(t)$ be a path from the identity to the swap gate in the unitary group on sites $j$ and $k$. We stack these gates together in a codimension 1 plane,
$\mathrm{SWAP}_{x,x+1}(t)=\bigotimes_{\vec{y}\in{\mathbb{Z}}^{d-1}}\mathrm{swap}_{(x,\vec{y}),(x+1,\vec{y})}(t).$ (6)
Then the path of QCA on a two-layer system
$u_{x,x+1}(v,t):=v_{x}\mathrm{SWAP}_{x,x+1}(t)v_{x}^{\dagger}\mathrm{SWAP}_{x,x+1}(t)^{\dagger}$ (7)
interpolates between the identity and $v_{x}v_{x+1}^{\dagger}$, where $v_{x}$ acts as $v$ on the $(d-1)$-dimensional surface $\{x\}\times{\mathbb{Z}}^{d-1}$. Finally, the swindle map is
$S(v,t)=\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}(v,t)\right)\left(\prod_{x\in 2{\mathbb{Z}}}u_{x,x+1}(v,t)\right),$ (8)
which is continuous in $t$ and $v$ and equal to the identity for $(t,v)\in\{0,1\}\times\mathrm{QCA}_{d-1}$ (Fig. 2).
Assuming that any loop of QCA can be restricted to open boundary conditions such that it remains a QCA, we show that $S$ is a homotopy equivalence in Appendix B. Formal details are deferred to Ref. [107].
Iteratively applying Eq. (5), we reduce the classification of driven localized phases in any dimension to either the classification of static QCA, or a driven classification of zero-dimensional QCA,
$\pi_{n}(\mathrm{QCA}_{d})\cong\left\{\begin{array}[]{l l}\pi_{0}(\mathrm{QCA}_{d-n})&\quad\text{for }d\geq n,\\ \pi_{n-d}(\mathrm{QCA}_{0})&\quad\text{for }d<n.\end{array}\right.$ (9)
The space $\mathrm{QCA}_{0}$ is a stabilization of the projective unitary group by stacking ancillae. We find its homotopy groups in Appendix C. Using the known classification of one- and two-dimensional QCA, we have all the homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ for spin systems with $d-n\leq 2$. These are listed in Tab. 1, and the corresponding classification of localized phases (excluding permutation QCA) is listed in Tab. 2. Comparing to the classification of invertible states shows that all the nontrivial phases in the tables are ALT phases.
$q$ $3$ $2$ $1$ 00 $-1$ $-2$ $-3$ $-4$ $-5$
$h^{q}$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$ 00 ${\mathbb{Q}}$
$h_{\mathrm{inv}}^{q}$ 00 ${\mathbb{Z}}$ 00 00 00 ${\mathbb{Z}}$ 00 00 00
Table 1: The homotopy groups $\pi_{n}(\mathrm{QCA}_{d})=:h^{q}$ are a function of $q=d-n$, as is $\pi_{n}(\mathrm{inv}_{d})=:h_{\mathrm{inv}}^{q}$. For spin systems $h^{3}$ is conjectured to be the Witt group of modular tensor categories [56, 108, 55, 109, 57]; $h^{2}$ is trivial [61]; $h^{1}$ is isomorphic to $\log{\mathbb{Q}}^{\times}$ [58]; and $h^{q}$ for $q\leq 0$ is trivial for even $q$, ${\mathbb{Q}}/{\mathbb{Z}}$ for $q=-1$, and ${\mathbb{Q}}$ for odd $q\leq-3$ (Appendix C). In all groups in the table, the group operation is addition.
$d\;\;$ $\;\;n$ 00 $1$ $2$ $3$ $4$ $5$
00 00 00 00 ${\mathbb{Q}}$ 00 ${\mathbb{Q}}$
$1$ 00 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$ 00
$2$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00 ${\mathbb{Q}}$
$3$ 00 $\log{\mathbb{Q}}^{\times}$ 00 ${\mathbb{Q}}/{\mathbb{Z}}$ 00
Table 2: The classification of strong invariants in $n$-tone driven $d$-dimensional localized phases of spins with short-range entanglement. The classifying group differs from $\pi_{n}(\mathrm{QCA}_{d})$ for $(d,n)\in\{(0,1),(1,0)\}$ due to the non-uniqueness of the parameterization by QCA. All nontrivial elements of this table correspond to ALT phases.
SALT phases.—Two systems belong to the same symmetric localized phase if they can be deformed into one another without breaking some symmetry or delocalizing. If the short-ranged entangled MBL Hamiltonian $H$ commutes with an on-site unitary abelian symmetry, then it can still have stable and complete LIOMs [67, 46, 54, 110], and the correspondence with QCA continues to hold. Thus, symmetry-enriched ALT phases (SALT phases, symmetric localized phases without eigenstate order) can be classified by symmetric QCA. (We do not consider weak symmetries, where the symmetry only exists on average [111, 112].) In fact, at the level of QCA, there is no need to restrict to abelian groups.
The conjectured classification of symmetry-enriched invertible states can be stated in terms of just the space $\mathrm{inv}_{d}$ and the abstract symmetry group $G$ [96, 66, 95, 102, 113, 103, 104]. However, the classification of symmetric QCA depends not just on $G$, but on the specific representation $\rho$ of the symmetry action [93, 114]. We compute the classification of symmetric QCA, and thus SALT phases, with $d-n\leq 1$ and any on-site unitary representation of any group in Appendix D. The classification of symmetric QCA closely mirrors Tab. 1, except that it includes additional subgroups in $d-n\geq 0$ [given by $H^{d-n+1}(G;U(1))$ for finite $G$] related to nontrivial SPT states or pumps thereof; moreover, rational numbers $p/q\in{\mathbb{Q}}$ in Tab. 1 are replaced with formal ratios of representations, $\rho/\sigma$. Addition and multiplication for $\rho/\sigma$ is defined through the same rules as for ratios of integers, but replacing “$+$” with $\oplus$” and “$\times$” with $\otimes$”. In an appropriate coincidence, this mathematical structure is known as an algebraic localization.
To define the algebraic localization precisely, we must specify the set of ancillae representations,
=$\mathcal{R}=\{\rho\in\mathrm{Rep}(G)\,:\,\text{some ancillae carry }\rho\},$ (10)
which is closed under tensor product—if $\rho_{1}$ and $\rho_{2}$ can be added then $\rho_{1}\otimes\rho_{2}$ corresponds to adding them both. The algebraic localization with denominators in $\mathcal{R}$ is the quotient
$\mathcal{R}^{-1}R_{G}=(R_{G}\times\mathcal{R})/\sim\quad\text{where }(\rho_{1},\sigma_{1})\sim(\rho_{2},\sigma_{2})\\ \Leftrightarrow\exists\rho\in\mathcal{R}\text{ such that }\rho\otimes\rho_{1}\otimes\sigma_{2}\simeq\rho\otimes\rho_{2}\otimes\sigma_{1},$ (11)
where $R_{G}$ is the representation ring of $G$, and $\simeq$” is unitary equivalence of representations. Equivalence classes of pairs $(\rho,\sigma)$ are denoted $\rho/\sigma$.
The representation theory of any group which is of relevance to physics is understood in extraordinary detail, so calculating $\mathcal{R}^{-1}R_{G}$, and thus the classification of low-dimensional SALT phases, becomes straightforward. For instance, the classification of $(d,n)=(2,1)$ $U(1)$-SALT phases in Ref. [26] by rational polynomials follows from observing that $R_{U(1)}\cong{\mathbb{Z}}[x,x^{-1}]$ is isomorphic to the ring of integer Laurent polynomials. The fact that including the regular representation, $\rho_{\text{reg}}$, for a finite group in $\mathcal{R}$ leads to a collapse in the classification [93] follows from the fact that $\rho_{\text{reg}}$ is a zero divisor in $R_{G}$. This means Eq. (11) can be satisfied even if $\rho_{1}\otimes\sigma_{2}\not\simeq\rho_{2}\otimes\sigma_{1}$.
Figure 3: Sketch of the pulse schedule for the models defined by the swindle map, with resulting unitary evolution $\tilde{S}$. Sites are coupled along even bonds (red, $t\in[0,\tfrac{1}{3})$) and then along odd bonds (blue, $t\in[\tfrac{1}{3},\tfrac{2}{3})$). Finally, an on-site disorder potential may be applied to promote localization (gray, $t\in[\tfrac{2}{3},1]$).
Models.—While the swindle map [Eq. (8), Fig. 2] was introduced as an abstract tool, it also defines a class of soluble models of $n$-tone-driven (S)ALT phases given any $(n-1)$-parameter family of QCA in one dimension lower. As these models are finite depth circuits, they are also well suited for realization in quantum simulator experiments [63, 64, 65] (Fig. 3).
As an example, Fig. 1(b) illustrates a three-dimensional phase with two driving tones which pumps anomalous Floquet phases to its boundary, and thus has a quantized surface magnetization [52, 27]. This is obtained by applying the swindle map to any model of the anomalous Floquet phase [22, 23].
As three-dimensional systems are typically still challenging for quantum simulators, in Appendix E we construct a soluble model of the quasiperiodic energy pump (QP pump) [31, 32, 53], a one-dimensional two-tone-driven SALT phase with $U(1)$ symmetry. The QP pump supports edge states which pump energy between the two driving fields. The classification of SALT phases shows that the QP pump is stable to strong interactions, provides a nearest-neighbor qubit model for this phase, and indicates that forgetting the $U(1)$ symmetry puts the model in the phase of Refs. [29, 30].
Discussion.—In the context of exactly soluble commuting models and driven phases of localized systems, authors have repeatedly expressed confusion regarding the position of their models within existing classifications [23, 29, 26, 91]. Most of these are classifications of states [66, 115, 28], while localized models (including any commuting projector model) impose nontrivial structure on the entire spectrum. This structure need not be captured by any single state, making such classifications of limited applicability. This point is implicit in the literature [23, 91, 31, 50], but the lack of a systematic and general study seems to have allowed confusion to persist.
Indeed, both the study of ground state phases and localized phases are rich mathematical problems with important implications for disparate physical systems. While the two problems are related (by $e_{0}^{*}$, with short-ranged entanglement), neither can be regarded as a special case of the other.
The general relationship between (S)ALT phases and QCA we have demonstrated, as well as the $\Omega$-spectrum property which relates static (S)ALT phases to driven phases, extends to fermionic systems essentially without modification (Appendix A-B[56]. The $\Omega$-spectrum structure of the classification can also be exploited to classify SALT phases with lattice symmetries [116, 103, 104]. The more challenging direction of generalization is to include topological order (long-ranged entanglement), where the correspondence with QCA breaks down [20]. This case is of particular interest given recent proposals of using driven topologically ordered localized models to perform quantum computation [117, 118, 119, 120]. The notion of measurement QCA introduced by Ref. [121] may provide a route towards classifying localized phases with topological order, which we will explore in future work.
The authors are particularly grateful to R. Kobayashi for collaboration during the early stages of this project. The authors also thank M. Barkeshli, A. Chandran, L. Fidkowski, S. Pace, R. Thorngren, Y.-T. Tu, B. Ware, and C. Zhang for useful discussions. This work is supported by the Laboratory for Physical Sciences (DML). Research at Perimeter Institute is supported in part by the Government of Canada through the Department of Innovation, Science and Economic Development, and by the Province of Ontario through the Ministry of Colleges and Universities (DVE).
Appendix A From ALT phases to QCA
Our primary tool for studying localized topological phases is a mapping from quantum cellular automata (QCA) to many-body localized (MBL) Hamiltonians with short-ranged entanglement. Explicitly, we define a locality preserving unitary which maps physical on-site degrees of freedom to the conserved local integrals of motion (LIOMs).
We formalize our definitions of “MBL Hamiltonian” and “short-ranged entanglement” thereof in Appendix A.1. The main point of this section and Appendix A.2 is that these physical notions reproduce the formal requirements of [56, Theorem II.4]. Thus, we realize abstract restrictions on commuting projector models as physical properties of localized systems. [56, Theorem II.4] then states that the local integrals of motion (LIOMs) of the MBL Hamiltonian can be prepared by a QCA. We make an equivalent proof in Appendix A.2. The QCA in question is not unique, and we characterize the redundancy in Appendix A.3. We also discuss localized phases of driven systems, which we also characterize (non-uniquely) by parameterized families of QCA.
While our discussion is phrased for spin systems, the results we review all hold for fermions as well [56].
A.1 Setup
While all classification results in this work are for spin systems on the integer lattice ${\mathbb{Z}}^{d}$, the correspondence between localized phases and QCA holds for either spins or fermions and any locally finite set of sites. That is, any subset of space with finite diameter contains finitely many degrees of freedom. We will call the of points in space supporting some degree of freedom a lattice, though it does not need to have any further regularity.
Formally, we consider a metric space $X$ and a lattice $K\subseteq X$ (which may include the same point multiple times) with associated finite dimensional local Hilbert spaces $\mathcal{H}_{k}$. This specializes our discussion to spin systems, though a parallel discussion holds for systems with fermions [56]. The total Hilbert space is the tensor product of the local Hilbert spaces,
$\mathcal{H}=\bigotimes_{k\in K}\mathcal{H}_{k}.$ (12)
(We will not need to be concerned with precise definitions of the infinite tensor product.)
A Hamiltonian on $X$ is a Hermitian operator $H$ on $\mathcal{H}$. We will consider fully localized Hamiltonians with strictly local LIOMs. The generic case is for the LIOMs to be quasilocal (with exponential tails), but for the purpose of mathematical proofs it is convenient to require that LIOMs have finite support on the lattice.
Definition A.1.
The Hamiltonian $H$ is range-$r$ fully many-body localized if there is a lattice of points $J\subseteq X$ (points may occur more than once) and Hermitian operators $\{\tau^{z}_{j}\}_{j\in J}$, called LIOMs, of the form
$\tau^{z}_{j}\in\mathbbm{1}\otimes\bigotimes_{k\in B_{r}(j)\cap K}\mathrm{End}(\mathcal{H}_{k}),$ (13)
[where $B_{r}(j)$ is the ball of radius $r$ around $j$] such that $[\tau^{z}_{j},H]=[\tau^{z}_{j},\tau^{z}_{j^{\prime}}]=0$. Further, the set of LIOMs must be maximal, such that a local operator $A$ commutes with all LIOMs only if $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$.
This definition should be compared to [56, Definition II.1].
We further say that $H$ is short-range entangled if all its excitations can be created and destroyed locally. This excludes models with nontrivial anyons, where there are local excitations which cannot be removed except by annihilating them with other, potentially distant, excitations. It also excludes symmetry breaking, where domain walls also cannot be destroyed locally. We formalize this notion with the following convenient definition.
Definition A.2.
The Hamiltonian $H$ is range-$r$ fully many-body localized with short-ranged entanglement if it is range-$r$ fully many-body localized and additionally there are unitary operators
$\tau^{x}_{j}\in\mathbbm{1}\otimes\bigotimes_{k\in B_{r}(j)\cap K}\mathrm{End}(\mathcal{H}_{k}),$ (14)
such that $[\tau^{x}_{j},\tau^{z}_{j^{\prime}}]=[\tau^{x}_{j},\tau^{x}_{j^{\prime}}]=0$ for $j\neq j^{\prime}$ and $\tau^{x}_{j}$ cyclically permutes the eigenspaces of $\tau^{z}_{j}$. That is, denoting the ordered set of distinct eigenvalues of $\tau^{z}_{j}$ as $(\lambda^{j}_{0},\ldots\lambda^{j}_{N_{j}-1})$ and the associated eigenspaces $\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})$, we have
$\tau^{x}_{j}\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})=\mathrm{eig}(\lambda^{j}_{l+1\bmod N_{j}},\tau^{z}_{j}).$ (15)
This definition should be compared to [56, Definition II.3]. One can also compare this notion of “short-ranged entanglement” for Hamiltonians to the usual notion of short-ranged entanglement for states. A state $|\psi\rangle$ may be said to be short-range entangled if, for any other state $|\psi^{\prime}\rangle$ which differs from $|\psi\rangle$ only in some ball $B_{r}(x)$ [$\langle\psi|A|\psi\rangle=\langle\psi^{\prime}|A|\psi^{\prime}\rangle$ for any local $A$ not supported in $B_{r}(x)$] there is a local unitary $\tau^{x}_{\psi\psi^{\prime}}$ supported in $B_{r}(x)$ such that $\tau^{x}_{\psi\psi^{\prime}}|\psi^{\prime}\rangle=|\psi\rangle$. Any eigenstate of an $H$ satisfying Definition A.2 is short-range entangled in this sense (the required $\tau^{x}_{\psi\psi^{\prime}}$ can be constructed from the $\tau^{x}_{j}$ operators), but the converse—that $\tau^{x}_{j}$ operators exist for any MBL Hamiltonian $H$ whose eigenstates are all short-range entangled—is not obvious.
Note that a commuting model is characterized less by the Hamiltonian $H$, and more by the algebra of the operators $\{\tau^{z}_{j},\tau^{x}_{j}\}_{j\in J}$. Indeed, focusing on LIOM operator algebras will be much more useful than the Hamiltonian itself in this context. In this way, we also avoid needing to be specific regarding the resolution of degeneracies in the Hamiltonian. A classification more directly based on the topology of the space of MBL Hamiltonians would need a definition of MBL which excludes highly degenerate Hamiltonians, such as the zero Hamiltonian, from being MBL. Otherwise all MBL Hamiltonians $H_{0}$ and $H_{1}$ would be connected by a continuous path with $H_{t}=(1-2t)H_{0}$ for $t\in[0,\tfrac{1}{2})$ and $H_{t}=(2t-1)H_{1}$ for $t\in[\tfrac{1}{2},1]$. Focusing on LIOMs captures the continuous change of local properties of the model while allowing for accidental degeneracies, so we will not address how to formulate a definition of MBL which avoids this issue.
For convenience, we define some abbreviated notation for operator algebras. For $S\subseteq X$ define
$\mathcal{O}_{S}=\mathbbm{1}\otimes\bigotimes_{k\in S\cap K}\mathrm{End}(\mathcal{H}_{k})\cong\bigotimes_{k\in S\cap K}\mathrm{Mat}_{M_{k}}({\mathbb{C}}),$ (16)
where $M_{k}=\mathrm{dim}(\mathcal{H}_{k})$ and $\mathrm{Mat}_{M_{k}}({\mathbb{C}})$ is the algebra of $M_{k}\times M_{k}$ matrices over the complex numbers. Similarly define the algebra generated by the $\tau$ operators within some region as
$\mathcal{T}_{S}=\langle\tau^{z}_{j},\tau^{x}_{j}\rangle_{j\in S\cap J}.$ (17)
Definition A.3.
A unitary operator $V$ on $\mathcal{H}$ is a range-$r$ quantum cellular automaton (QCA) if
$V\mathcal{O}_{S}V^{\dagger}\subseteq\mathcal{O}_{B_{r}(S)}$ (18)
for all $S\subseteq X$.
Here, $B_{r}(S)$ is the ball of radius $r$ around the set $S$. Properly, a QCA should be a projective unitary, and this distinction will be important in later appendices, but not here.
Reference [51] defines a commuting model to be “without intrinsic topological order” when the LIOMs $\tau^{z}_{j}$ can be prepared from single site operators
$\sigma^{z}_{k}\in\mathbbm{1}\otimes\mathrm{End}(\mathcal{H}_{k})$ (19)
by a QCA $V$. Any such Hamiltonian also clearly has short-ranged entanglement by our Definition A.2, as we can construct $\tau_{k}^{x}=V\sigma^{x}_{k}V^{\dagger}$ from a matrix $\sigma^{x}_{k}\in\mathrm{End}(\mathcal{H}_{k})$ which acts as a cyclic permutation on the eigenvectors of $\sigma^{z}_{k}$. The result of Appendix A.2 is that the QCA $V$ always exists given Definition A.2 as well, and as such the two notions short-ranged entanglement and lack of intrinsic topological order are equivalent.
A.2 Existence
In this appendix, we prove the following theorem.
Theorem A.1.
If $H$ is range-$r$ many-body localized with short-ranged entanglement, then there is a range-$2r$ QCA $V$ such that each $V^{\dagger}\tau^{z}_{j}V$ is a sum of products of commuting single-site operators $\sigma^{z}_{k}$.
In fact, after a redefinition of LIOMs to another equivalent set, we will have that $V^{\dagger}\tau^{z}_{j}V$ is a single site operator.
The proof is essentially the same as [56, Theorem II.4]. The main insight here is that the abstract requirements of that theorem correspond to natural physical requirements for an MBL Hamiltonian.
The strategy will be to show that each of the algebras
$\mathcal{T}_{j}=\langle\tau^{z}_{j},\tau^{x}_{j}\rangle$ (20)
is isomorphic to a matrix algebra. This implies that there is an isomorphism between $\mathcal{T}_{j}$ and $\mathcal{O}_{k}$ whenever they have the same dimension. [All isomorphisms will be *-isomorphisms—that is an isomorphism $\alpha$ such that $\alpha(A^{\dagger})=\alpha(A)^{\dagger}$ for all operators $A$. All such automorphisms of this kind between matrix algebras can be expressed as a conjugation by a unitary.] To find a QCA $V$ which implements this isomorphism simultaneously for all sites, we need to match the algebras $\mathcal{O}_{k}$ to $\mathcal{T}_{j(k)}$ locally and bijectively. This can be phrased as a graph problem, and Hall’s marriage theorem [122, 123] implies that this problem always has a solution.
Lemma A.2.
There is an isomorphism
$\mathcal{T}_{j}\cong\mathrm{Mat}_{N_{j}}({\mathbb{C}})$ (21)
for each $j\in J$.
Proof.
Wedderburn’s classification of finite dimensional semisimple algebras implies that $\mathcal{T}_{j}$ is isomorphic to a full matrix algebra if and only if it is central. That is, $A\in\mathcal{T}_{j}$ commutes with all of $\mathcal{T}_{j}$ only if $A\in{\mathbb{C}}\mathbbm{1}.$
Suppose $A\in Z(\mathcal{T}_{j})$ is in the center of $\mathcal{T}_{j}$. Then it must commute with all the LIOMs—it commutes with $\tau^{z}_{j}$ by assumption and it commutes with the other $\tau^{z}_{l}$ because it is contained in $\mathcal{T}_{j}$. Then, by Definition A.1, $A$ is a sum of products of LIOMs. However, as $A\in\mathcal{T}_{j}$, it must be a polynomial in $\tau^{z}_{j}$. In particular,
$A=\bigoplus_{l=1}^{N_{j}-1}a_{l}\mathbbm{1}_{\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})}$ (22)
for some $a_{l}\in{\mathbb{C}}$. Evaluating the commutator with $\tau^{x}_{j}$, we have
$\tau^{x\dagger}_{j}A\tau^{x}_{j}=\bigoplus_{l=1}^{N_{j}-1}a_{l+1\bmod N_{j}}\mathbbm{1}_{\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})}=A,$ (23)
so all the $a_{l}=a$ are equal and $A\in{\mathbb{C}}\mathbbm{1}$. ∎
It is convenient to require that all the $N_{j}$ and $M_{k}=\mathrm{dim}(\mathcal{H}_{k})$ are prime. This can always be achieved by redefining the lattices $J$ and $K$. Indeed, focusing on $N_{j}$, we can make a prime factorization
$N_{j}=p_{0}p_{1}\cdots p_{m_{j}-1}$ (24)
(not all necessarily distinct) and express
$\mathcal{T}_{j}\cong\mathrm{Mat}_{N_{j}}({\mathbb{C}})\cong\bigotimes_{n=0}^{m_{j}-1}\mathrm{Mat}_{p_{n}}({\mathbb{C}}).$ (25)
Then we have that $\mathcal{T}_{j}=\prod_{n=0}^{m-1}\mathcal{T}_{j_{n}}$ with each tensor factor being of prime dimension. Further, we can make this factorization such that the eigenbasis of $\tau^{z}_{j}$ becomes a product basis. Indeed, we have that any $l\in\{0,\ldots,N_{j}-1\}$ can be uniquely expressed as a sum of integers
$l=\sum_{n=0}^{m_{j}-1}\mu^{l}_{n}\prod_{q=0}^{n-1}p_{q}$ (26)
with $\mu^{l}_{n}\in\{0,\ldots,p_{n}-1\}$. (Empty sums are taken to be 0 and empty products are 1.) Then we can define Hermitian projectors $\pi^{j}_{\mu n}$ (with $\mu\in\{0,\ldots,p_{n}-1\}$ and $n\in\{0,\ldots,m_{j}-1\}$) onto the direct sum of eigenspaces $\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})$ such that $\mu^{l}_{n}=\mu$. Note that all $\pi^{j}_{\mu n}$ with the same $n$ have the same rank, specifically $N_{j}/p_{n}$. As $\mathcal{T}_{j}$ is a full matrix algebra, we can find $\tau^{x}_{j_{n}}$ such that
$\tau^{x}_{j_{n}}\pi^{j}_{\mu n}\tau^{x\dagger}_{j_{n}}=\pi^{j}_{(\mu+1\bmod p_{n})n}$ (27)
and such that $\tau^{x}_{j_{n}}$ commutes with all the other projectors. Indeed, $\tau^{x}_{j_{n}}$ is defined by cyclically permuting the value of $\mu^{l}_{n}$ in the decomposition of $l$ for $\mathrm{eig}(\lambda^{j}_{l},\tau^{z}_{j})$. By essentially the same calculation as Lemma A.2, $\langle\tau^{x}_{j_{n}},\pi^{j}_{\mu n}\rangle_{\mu=0}^{p_{n}-1}$ is isomorphic to a full matrix algebra, and so occurs as a tensor factor in $\mathcal{T}_{j}$. Then we make the decomposition of $\mathcal{T}_{j}$ as
$\mathcal{T}_{j}\cong\bigotimes_{n=0}^{m_{j}-1}\langle\tau^{x}_{j_{n}},\pi^{j}_{\mu n}\rangle_{\mu=0}^{p_{n}-1}.$ (28)
We pick new LIOMs $\tau^{z}_{j_{n}}=\sum_{\mu=0}^{p_{n}-1}\lambda^{j_{n}}_{\mu}\pi^{j}_{\mu n}$ with $\lambda^{j_{n}}_{\mu}\in{\mathbb{R}}$ all distinct. These commute with $\tau^{z}_{j}$, and so are each actually functions of $\tau^{z}_{j}$ by Definition A.1. On the other hand, they generate all diagonal operators in $\mathcal{T}_{j}$ in the $\tau^{z}_{j}$ eigenbasis, so $\tau^{z}_{j}$ can also be recovered from these new LIOMs. Thus, the new LIOMs are equivalent to the old LIOM.
Replace the lattice $J$ with a new lattice $J^{\prime}$ which consists of the old lattice sites $j$ each repeated $m_{j}$ times, and associated LIOMs $\tau^{z}_{j_{n}}$ and algebras $\mathcal{T}_{j_{n}}$. This lattice has the properties we need. We henceforth drop primes, and just use $J$. Similarly, we can ensure that each of the local Hilbert spaces $\mathcal{H}_{k}$ is of prime dimension.
As $\mathcal{T}_{j}$ and $\mathcal{O}_{k}$ are both full matrix algebras, there is an isomorphism between them whenever they have the same dimension. To build these individual isomorphisms into a QCA, we need to make a matching between $\mathcal{T}_{j}$ and $\mathcal{O}_{k}$ algebras which are supported near one another. Define the bipartite graph $B=((J,K),E)$ with edges
$(j,k)\in E\iff N_{j}=M_{k}\text{ and }\mathrm{dist}(j,k)\leq r.$ (29)
Then we are looking for a perfect matching in this graph. That is, a bijection $k\mapsto j(k)$ such that $(j(k),k)\in E$ for all $k\in K$. Hall’s marriage theorem [122] (specifically, the infinite version due to Hall [123]—no relation) gives necessary and sufficient conditions for a perfect matching to exist. Define the neighborhood of $S\subseteq B$ as
$\partial S=\{b\in B\,:\,(s,b)\in E\text{ for some }s\in S\}.$ (30)
Then the marriage theorem states that if: (1) every vertex has finite degree, (2) for every finite $S\subseteq J$ we have
$|S|\leq|\partial S|,$ (31)
and similarly (3) for each finite $S\subseteq K$ we have $|S|\leq|\partial S|$, then the required perfect matching bijection exists. The criterion (31) is called the marriage condition.
Lemma A.3.
The graph $B$ has a perfect matching.
Proof.
We verify the marriage condition, Eq. (31).
Note that every vertex in $B$ has finite degree by the assumption that $K$ and $J$ are locally finite. Further observe that the graph $B$ splits into distinct connected components labelled by primes $p$ such that $N_{j}=M_{k}=p$. Thus, we can consider each component separately. Henceforth, we assume that we are working in a given component.
Let $S\subseteq J$ be a finite subset of LIOM sites. Then we have that $\mathcal{T}_{S}$ is a subalgebra in $\mathcal{O}_{B_{r}(S)}$. As both $\mathcal{T}_{S}$ and $\mathcal{O}_{B_{r}(S)}$ are full matrix algebras, $\mathcal{O}_{B_{r}(S)}$ must be a tensor product of $\mathcal{T}_{S}$ and its commutant in $\mathcal{O}_{B_{r}(S)}$. Comparing dimensions then shows that $\prod_{j\in S}N_{j}$ divides $\prod_{k\in B_{r}(S)}M_{k}$. Then, all of the primes $N_{j}=p$ must occur among the primes $M_{k}$, repeated according to their multiplicity. The sites in $B_{r}(S)$ with local Hilbert space dimension $M_{k}=p$ are precisely the elements of $\partial S$ in this component. Thus, we conclude $|S|\leq|\partial S|$.
In the other direction, let $S\subseteq K$ be a finite subset of lattice sites. Then we also have that $\mathcal{O}_{S}$ is a subalgebra of $\mathcal{T}_{B_{r}(S)}$, whereupon the proof proceeds as in the previous paragraph. Indeed, suppose that $A\in\mathcal{O}_{S}\subseteq\mathcal{O}_{B_{2r}(S)}$. As $\mathcal{T}_{B_{r}(S)}\subseteq\mathcal{O}_{B_{2r}(S)}$ occurs as a tensor factor in $\mathcal{O}_{B_{2r}(S)}$, $A$ is a sum of products of operators in $\mathcal{T}_{B_{r}(S)}$ and its commutant. We need to show that if $A$ is in the commutant, then it is a scalar. Indeed, as their supports do not overlap, $A$ must also be in the commutant of $\mathcal{T}_{K\setminus B_{r}(S)}$. Thus, $A$ commutes with all of the LIOMs, and by Definition A.1 we must have $A\in\langle\tau^{z}_{j}\rangle_{j\in J}$. The only such operator which also commutes with all the $\tau^{x}_{j}$ operators is a scalar, $A=a\mathbbm{1}$. ∎
We can now prove Theorem A.1.
Proof of Theorem A.1.
Let $j:K\to J$ be a perfect matching as in Lemma A.3. By Lemma A.2 there is an isomorphism of algebras
$\alpha_{k}:\mathcal{O}_{k}\to\mathcal{T}_{j(k)}.$ (32)
Then
$\alpha=\bigotimes_{k\in K}\alpha_{k}$ (33)
is an algebra isomorphism between $\bigotimes_{k\in K}\mathcal{O}_{k}$ and $\bigotimes_{j\in J}\mathcal{T}_{j}$. It is bijective because the matching $k(j)$ is perfect, and it is a homomorphism because all the algebras $\mathcal{T}_{j}$ commute with one another.
Further, $\alpha$ is $2r$-locality preserving: we have $\mathrm{dist}(j(k),k)\leq r$ by the definition of the graph $B$, and then $\alpha(\mathcal{O}_{k})=\mathcal{T}_{j(k)}\subseteq\mathcal{O}_{B_{2r}(k)}$ by the fact that $\mathcal{T}_{j(k)}$ is supported in $\mathcal{O}_{B_{r}(j(k))}$ and the triangle inequality.
All such complex matrix algebra homomorphisms are of the form $\alpha(A)=VAV^{\dagger}$ for some unitary $V$, which is the QCA we need. It maps all LIOMs to single-site operators, thus meeting the requirements of the theorem. ∎
Define the single-site operators $\sigma^{z}_{k}=V^{\dagger}\tau^{z}_{j(k)}V$. Then we have that
$V^{\dagger}HV=\sum_{k}h_{k}\sigma^{z}_{k}+\sum_{kl}h_{kl}\sigma^{z}_{k}\sigma^{z}_{l}+\cdots$ (34)
is a sum of products of the commuting single-site operators $\sigma^{z}_{k}$ (by Definition A.1). We note that this is just a diagonalization of $H$ in the simultaneous eigenbasis of all the $\sigma^{z}_{j}$. The nontrivial conclusion of this section is that this diagonalization can be made locally.
This characterization of localized system with short-ranged entanglement by QCA motivates a similar definition for driven systems. We consider an $n$-tone-driven system to be localized with short-ranged entanglement when it has a generalized Floquet decomposition [28, 99]
$U(t)=V({\vec{\theta}}_{t})e^{-itH_{F}}V({\vec{\theta}}_{0})^{\dagger},$ (35)
where the static Floquet Hamiltonian
$H_{F}=\sum_{k}h_{k}\sigma^{z}_{k}+\sum_{kl}h_{kl}\sigma^{z}_{k}\sigma^{z}_{l}+\cdots$ (36)
is a sum of products of the fixed $\sigma^{z}_{k}$ operators and the micromotion operator $V({\vec{\theta}})$ is a continuous parameterized family of QCA. By a similar proof to the above, we see that this definition is equivalent to the existence of a maximal set of LIOMs $\tau^{z}_{j}({\vec{\theta}})$ with explicit continuous ${\vec{\theta}}$ dependence and a family of conjugate $\tau^{x}_{j}({\vec{\theta}})$ operators also with continuous ${\vec{\theta}}$ dependence. Alternatively, the definition is also equivalent to the quasienergy operator [124, 98, 125, 126, 127] having a complete set of LIOMs [99] and conjugate operators which create or destroy excitations locally.
The requirement of the family $\tau^{x}_{j}({\vec{\theta}})$ being continuous is nontrivial. For instance, even for 2-tone-driven qubits we may construct continuous families of LIOMs such that each family of LIOM eigenstates has a nonzero Chern number [127, 128]. However, there is an obstruction to continuously defining a family of conjugate operators, as the eigenstates cannot be given a smooth gauge. We do not consider such systems to be localized as they have continuous spectrum [128], and so do not enjoy the same stability to generic perturbations we expect of localized systems [99]. Indeed, in a frequency lattice picture, such systems are delocalized [128].
We emphasize that Floquet localized phases with short-ranged entanglement correspond to loops of QCA—not loops in the space of QCA modulo generalized permutations, even if QCA which differ by generalized permutations correspond to the same LIOMs. Indeed, a loop of QCA modulo permutations could correspond to a path from the identity QCA to a permutation. This path of QCA does not define a legitimate micromotion operator $V({\vec{\theta}})$.
A.3 Non-uniqueness
A QCA which maps $H$ to a sum of product of $\sigma^{z}_{k}$ operators always exists, but this QCA is not unique.
Indeed, the QCA $V$ is a unitary operator which diagonalizes $H$ in the simultaneous eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$. Even if $H$ is nondegenerate, such unitaries are only unique up to reordering of eigenvalues and multiplying eigenstates by arbitrary phases. Specifically, if $P$ is a unitary generalized permutation matrix (a unitary matrix with exactly one unit modulus complex number in each row and column) in the eigenbasis of $\{\sigma^{z}_{k}\}_{k\in K}$, then $(VP)^{\dagger}H(VP)$ is still diagonal. If $H$ is nondegenerate, the full redundancy of the correspondence between localized models with short-ranged entanglement and QCA is found by restricting $P$ to be a QCA.
Translations form an important class of QCA permutation operators [58, 61]. In one dimension, all QCA are given by translation operators, up to a finite depth circuit. Thus, while one-dimensional QCA are classified by $\log{\mathbb{Q}}^{\times}$ [58], all these nontrivial QCA land within the redundancy of the description of localized phases. As such, all one-dimensional localized phases of spins without symmetry are trivial. The same is not true of fermions, where the Kitaev chain [129] represents a nontrivial localized phase which can be prepared by a QCA [117]. It has invertible eigenstate order, and so is not an ALT phase.
Mathematically, the classification of localized phases with short-ranged entanglement is given by a quotient of $\pi_{0}(\mathrm{QCA}_{d})$. Denoting the set of generalized permutation QCA by $\mathrm{GP}_{d}$, there is a canonical inclusion map $\iota:\mathrm{GP}_{d}\to\mathrm{QCA}_{d}$, and an induced homomorphism $\iota^{*}:\pi_{0}(\mathrm{GP}_{d})\to\pi_{0}(\mathrm{QCA}_{d})$. We write the image of this homomorphism as $I_{0}=\iota^{*}(\pi_{0}(\mathrm{GP}_{d}))$. Then the quotient which classifies localized topological phases is
$\pi_{0}(\mathrm{QCA}_{d})/I_{0}.$ (37)
This group is isomorphic to $\pi_{0}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})$. Indeed, there is a fibration sequence
[…]
(where $q$ is the quotient map which sends a QCA $V$ to its coset $V\mathrm{GP}_{d}$) with associated long exact sequence of homotopy groups
$\cdots\to\pi_{0}(\mathrm{GP}_{d})\xrightarrow{\iota^{*}}\pi_{0}(\mathrm{QCA}_{d})\xrightarrow{q^{*}}\pi_{0}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})\to 0,$ (39)
and $\pi_{0}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})\cong\pi_{0}(\mathrm{QCA}_{d})/I_{0}$ follows from the first isomorphism theorem applied to $q^{*}$.
In the driven case, the redundancy is captured by a family of generalized permutation matrices $P({\vec{\theta}})$ with elements
$P_{\alpha\beta}({\vec{\theta}})=\delta_{\pi(\alpha)\beta}e^{-i{\vec{m}}_{\beta}\cdot{\vec{\theta}}-i\phi_{\beta}},$ (40)
where $\pi$ is a permutation of the simultaneous eigenstates of the $\{\sigma^{z}_{k}\}_{k\in K}$ operators and ${\vec{m}}_{\alpha}\in{\mathbb{Z}}^{n}$ are vectors of integers. [Such $P({\vec{\theta}})$ are generalized permutation matrices in a frequency lattice picture [31, 99].] Any $P$ of this form defines an identical unitary evolution (35) with
$V^{\prime}({\vec{\theta}})$ $=V({\vec{\theta}})P({\vec{\theta}}),$ (41)
$H_{F}^{\prime}$ $=H_{F}-i[\partial_{t}P({\vec{\theta}}_{t})]P({\vec{\theta}}_{0})^{\dagger}.$ (42)
Note that $i[\partial_{t}P({\vec{\theta}}_{t})]P({\vec{\theta}}_{t})^{\dagger}$ is a static diagonal operator with elements
$(i[\partial_{t}P({\vec{\theta}}_{t})]P({\vec{\theta}}_{t})^{\dagger})_{\alpha\beta}={\vec{m}}_{\pi^{-1}(\alpha)}\cdot{\vec{\omega}}\,\delta_{\alpha\beta},$ (43)
where ${\vec{\omega}}=(\omega_{1},\ldots,\omega_{n})$ is the vector of drive frequencies. If $P({\vec{\theta}})$ is also a QCA for all ${\vec{\theta}}$, then $V^{\prime}({\vec{\theta}})$ also defines an acceptable generalized Floquet decomposition.
The only example for which this redundancy affects the classification of localized phases for $d\leq 3$ is the classification of Floquet ($n=1$) phases in zero dimensions [31]. Looking ahead to Appendix C.2, we have that $\pi_{1}(\mathrm{QCA}_{0})={\mathbb{Q}}/{\mathbb{Z}}$, and that a complete index is given by the winding number $\tilde{w}_{1}[V]$. In zero dimensions, all unitaries are QCA, so the diagonal unitary with elements
$P_{m}(\theta)_{\alpha\beta}=e^{im\theta\delta_{\alpha 0}}\delta_{\alpha\beta}$ (44)
defines an acceptable $V^{\prime}(\theta)=V(\theta)P_{m}(\theta)$. However, we can compute that
$\tilde{w}_{1}[VP_{m}]=\tilde{w}_{1}[V]+\frac{m}{N}\mod 1,$ (45)
where $N$ is the Hilbert space dimension. Thus, multiplication by $P_{m}(\theta)$ can take $\tilde{w}_{1}[V^{\prime}]$ to any allowed value, and we conclude that all loops of zero dimensional QCA (loops of projective unitaries) define trivial localized phases.
More systematically, we can again classify driven localized phases by a quotient of the group $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$. The generalized permutations $P({\vec{\theta}})$ are particular maps from ${\mathbb{T}}^{n}$ to $\mathrm{QCA}_{d}$, and we denote the subgroup of $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]$ induced by the inclusion of the generalized permutations in the full space of maps from ${\mathbb{T}}^{n}$ to $\mathrm{QCA}_{d}$ by $I_{n}$. Then the classification of $n$-tone driven localized phases in $d$-dimensions with short-ranged entanglement is given by
$\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}\right]/I_{n}.$ (46)
This group is not isomorphic to $\pi_{0}\left[{\mathbb{T}}^{n},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]$ in general. Taking, for instance, $n=1$, there are loops in $\pi_{0}\left[{\mathbb{T}}^{1},\mathrm{QCA}_{d}/\mathrm{GP}_{d}\right]\cong\pi_{1}(\mathrm{QCA}_{d}/\mathrm{GP}_{d})$ which correspond to paths from the identity to a locally generated permutation. This is not a legitimate micromotion operator $V(\theta)$, as it is not periodic as a (projective) unitary.
Appendix B The $\Omega$-spectrum property
In topology, an $\Omega$-spectrum refers to a sequence of topological spaces $(Y_{d})_{d\in{\mathbb{Z}}}$ together with (weak) homotopy equivalences
$Y_{d-1}\simeq\Omega Y_{d}$ (47)
where $\Omega Y_{d}$ is the space of loops in $Y_{d}$ which begin and end at a specified base point. $\Omega$-spectra play a special role in topology, as they serve to define generalized cohomology theories—functors from the category of topological spaces to abelian groups obeying certain axioms [95, 102, 104, 103]. The functor in question is
$h^{d}_{Y}(X)=\pi_{0}\left[X,Y_{d}\right].$ (48)
The $\Omega$-spectrum property comes with substantial structure, and highly constrains the properties of the spaces $Y_{d}$.
We conjecture that the spaces of quantum cellular automata on the integer lattice ${\mathbb{Z}}^{d}$ stabilized by adding ancillae, $\mathrm{QCA}_{d}$, form an $\Omega$-spectrum,
$\mathrm{QCA}_{d-1}\simeq\Omega\mathrm{QCA}_{d}.$ (49)
For $d<0$, we define $\mathrm{QCA}_{d}=\Omega^{-d}\mathrm{QCA}_{0}$, making the negative degrees of this statement trivial. In positive degree, we propose that the swindle map $S$ [Eq. (8)] provides the homotopy equivalence. Thus, we propose there is a generalized cohomology theory
$h^{d}(X)=\pi_{0}\left[X,\mathrm{QCA}_{d}\right].$ (50)
In this appendix, we provide arguments in support of this conjecture. In Appendix B.1, we recall that if loops of QCA can be put on open boundary conditions such that they act by QCA on their edges—known as blending with the identity—then loops of QCA can all be parameterized as finite depth circuits. We also argue that this condition holds for loops of one-dimensional QCA. Then, in Appendix B.2, we construct an explicit deformation retract from the space of circuits to the image of the swindle map. As the swindle map is injective, this shows
$\mathrm{QCA}_{d-1}\cong\mathrm{im}(S)\simeq\Omega\mathrm{QCA}_{d}$ (51)
under the assumptions of Appendix B.1.
Making these arguments rigorous requires being more specific about the topology on $\mathrm{QCA}_{d}$, even in $d=1$. This can be done, but we defer such technicalities to Ref. [107].
While it may be useful to keep spin systems in mind, the arguments we present in this appendix are agnostic to local operators being fermionic or spin operators. Provided that the additional ancillae used to stabilize a system may support fermionic operators, everything in the following sections also applies to systems with fermions.
B.1 Blending and circuits
A blending between $\alpha$ and $\beta\in\mathrm{QCA}_{d}$ is a third QCA $\gamma$ such that
$\gamma(A)$ $=\alpha(A)\quad\text{for }\mathrm{supp}(A)\subseteq(-\infty,x_{l}]\times{\mathbb{Z}}^{d-1},$ (52)
$\gamma(A)$ $=\beta(A)\quad\text{for }\mathrm{supp}(A)\subseteq[x_{r},\infty)\times{\mathbb{Z}}^{d-1},$ (53)
where $\mathrm{supp}(A)$ is the smallest subset of ${\mathbb{Z}}^{d}$ on which the operator $A$ acts non-trivially, and $x_{l,r}$ are positions along the first axis in ${\mathbb{Z}}^{d}$. That is, $\gamma$ agrees with $\alpha$ to the left of some cut, and agrees with $\beta$ to the right of the cut. We can similarly define a blending $\gamma({\vec{\theta}})$ between parameterized families of QCA $\alpha({\vec{\theta}})$ and $\beta({\vec{\theta}})$. Two (families of) QCA are said to blend into each other if a blending exists between them.
Two QCA blend into one another if and only if they are related by a finite depth circuit [94]. That is, $\alpha$ blends with $\beta$ at any position of the cut $x$ if and only if $\alpha\beta^{-1}$ is equal to a finite depth unitary circuit. (Actually, it is sufficient for a blending to exist at any position [109].)
Indeed, it is sufficient to consider a blending between $\gamma=\alpha\beta^{-1}$ and the identity. Then $\gamma$ should be equal to a finite depth circuit. Let the blending at the cut across $\{x\}\times{\mathbb{Z}}^{d-1}$ be $\tilde{\gamma}_{x}$, and coarse grain such that $\tilde{\gamma}$ agrees with $\gamma$ for $x^{\prime}<x$ and acts as the identity for $x^{\prime}>x$. Then we have that $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ and $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$ are QCA supported on $\{x,x+1\}\times{\mathbb{Z}}^{d-1}$.
We can split $\gamma$ into a depth two circuit by iteratively splitting off gates of the form $\tilde{\gamma}_{x}\tilde{\gamma}_{x+1}^{-1}$ or $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x+1}$. (The gates are themselves QCA in $d-1$ dimensions.) First, observe that we can split $\gamma$ into a product of two semi-infinite QCA
$\gamma=\tilde{\gamma}_{0}(\tilde{\gamma}_{0}^{-1}\gamma).$ (54)
Then we can express the $\tilde{\gamma}_{0}$ factor as
$\tilde{\gamma}_{0}=(\tilde{\gamma}_{0}\tilde{\gamma}_{-1}^{-1})\tilde{\gamma}_{-1}.$ (55)
The first factor is the gate we need. We can proceed similarly with the second factor, decomposing $\tilde{\gamma}_{-1}=\tilde{\gamma}_{-2}(\tilde{\gamma}^{-1}_{-2}\tilde{\gamma}_{-1})$ and so on, alternating whether we perform the multiplication by $\tilde{\gamma}^{-1}_{x}\tilde{\gamma}_{x}$ on the right or $\tilde{\gamma}_{x}\tilde{\gamma}^{-1}_{x}$ on the left. Mirroring the same procedure for positive $x$, we obtain a depth-two circuit decomposition
$\gamma=\left(\prod_{x\in 2{\mathbb{Z}}+1}w_{x,x+1}\right)\left(\prod_{x\in 2{\mathbb{Z}}}w_{x,x+1}\right).$ (56)
The converse direction for the equivalence of blending and circuit equivalence is straightforward. If $\gamma$ is a finite depth circuit, we can simply delete all gates to one side of any cut to obtain a blending with the identity.
It is intuitive that a blending should exist for any path of QCA beginning at the identity. Such a path of unitary operators $\gamma(t)$ can be generated by a local Hamiltonian $H(t)$ if it is differentiable. Then tracing over the sites to the right of any cut gives a new Hamiltonian $\mathrm{Tr}_{(x,\infty)\times{\mathbb{Z}}^{d-1}}{H(t)}$ which, assuming it continues to generate a QCA, generates a blending between $\gamma(t)$ and the identity. This is certainly true in one dimension, where any unitary evolution at the zero-dimensional edge is a QCA, but in higher dimensions the new Hamiltonian evolution may develop exponential tails along the edge. This is not much of a problem physically, but such tails are an obstacle to performing proofs.
Nonetheless, we expect that the path-space $P\mathrm{QCA}_{d}$ can be continuously parameterized in terms of a strictly local path of circuits in dimensions higher than one too. The loop space $\Omega\mathrm{QCA}_{d}$ can then be parameterized in terms of circuits which begin and end at the identity. Note that this does not require that the individual gates must return to the identity, only that they cancel at the end of the path.
In Ref. [107] we study this structure in greater detail, and characterize the circuit liftings implicit here as a fibration on the space of QCA.
B.2 Deformation to the swindle
We now argue that Eq. (51) holds.
First, we show that $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is injective, such that $\mathrm{im}(S)\cong\mathrm{QCA}_{d-1}$ is homeomorphic to $\mathrm{QCA}_{d-1}$. That is, $S(v,t)=S(w,t)$ for all $t$ only if $v=w$. Indeed, multiplying $S(v,t)$ on the left by the inverse of the top layer in $S(w,t)$ and on the right by the inverse of the bottom layer (Fig. 2), we have
$\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}(w,t)^{\dagger}u_{x,x+1}(v,t)\right)\\ \times\left(\prod_{x\in 2{\mathbb{Z}}}u_{x,x+1}(v,t)u_{x,x+1}(w,t)^{\dagger}\right)=\mathbbm{1}.$ (57)
This equation can only be satisfied if, for all $t\in[0,1]$,
$u_{2x,2x+1}(v,t)u_{2x,2x+1}(w,t)^{\dagger}$ $=q_{2x}(t)q_{2x+1}(t)$ (58)
$u_{2x-1,2x}(w,t)^{\dagger}u_{2x-1,2x}(v,t)$ $=q_{2x-1}(t)^{\dagger}q_{2x}(t)^{\dagger}$ (59)
for some sequence of $q_{x}(t)$ supported on $\{x\}\times{\mathbb{Z}}^{d-1}$. However, the gates $u_{x,x+1}$ are translationally invariant, so considering $x=0$ gives
$u_{0,1}(v,t)u_{0,1}(w,t)^{\dagger}=q_{0}(t)q_{1}(t)=u_{0,1}(v,t)^{\dagger}u_{0,1}(w,t),$ (60)
where the second equality comes from translating Eq. (59) by one unit towards positive $x$ and inverting the equation. We see that $u_{0,1}(v,t)^{2}=u_{0,1}(w,t)^{2}$. Making a spectral decomposition of both sides, we conclude that $u_{0,1}(v,t)=u_{0,1}(w,t)$ up to multiplying eigenvalues by ±$\pm 1$. However, these paths both start at $t=0$ with the same eigenvalues $\lambda=+1$, which remain nonzero throughout all $t$. Thus, all the discrete choices of ±$\pm 1$ must be $+1$, and $u_{0,1}(v,t)=u_{0,1}(w,t)$. Then comparing these at $t=1$ gives $v_{0}v^{\dagger}_{1}=w_{0}w^{\dagger}_{1}$, and hence $v=w$ up to a phase. We regard QCA as projective unitaries, so $v=w$.
We now assume that all loops of QCA $\gamma(t)$ can be represented by paths of circuits (Appendix B.1)
$\gamma(t)=\left(\prod_{x\in 2{\mathbb{Z}}+1}w_{x,x+1}(t)\right)\left(\prod_{x\in 2{\mathbb{Z}}}w_{x,x+1}(t)\right)$ (61)
such that $w_{x,x+1}(0)=\mathbbm{1}$ and the circuit is the identity at $t=1$. This requires that $w_{x,x+1}(1)=v^{\gamma}_{x}\otimes v^{\gamma\dagger}_{x+1}$ for some sequence $v^{\gamma}_{x}\in\mathrm{QCA}_{d-1}$ which are not necessarily all equal.
Figure 4: The loop of finite depth circuits $\gamma(t)\otimes S(v_{0}^{\gamma},s)^{\dagger}$ is deformed to the constant loop at the identity as a function of $s\in[0,1]$. (a) At $s=0$, the circuit consists of two layers. Between $s=0$ and $s=1/2$ the gates $u^{\dagger}_{x,x+1}(v_{0}^{\gamma},t)$ in the top layer (red) are swapped to the bottom layer while having their order of multiplication reversed. (b) The gate $u^{\dagger}_{0,1}(v_{0}^{\gamma},t)w_{0,1}(t)$ [$u^{\dagger}_{-1,0}(v_{0}^{\gamma},t)w_{-1,0}(t)$] acts as the identity on site 00 for $t\in\{0,1\}$. It can thus be swapped (over $s\in[\tfrac{1}{2},\tfrac{3}{4}]$) so that it acts on site $1$ [$-1$] and its ancilla while ensuring that the whole circuit is still a loop. (c) Between $s=\tfrac{3}{4}$ and $s=1$ the circuit (red) is “rolled up”, such that it acts as the identity (gray dots) in any finite region.
Then, up to stable equivalence, we have
$\gamma(t)=\gamma(t)\otimes\mathbbm{1}=\gamma(t)\otimes S(v^{\gamma}_{0},t)^{\dagger}S(v^{\gamma}_{0},t).$ (62)
We now make a sequence of deformations which cancel the $\gamma(t)$ and $S(v_{0},t)^{\dagger}$ parts of Eq. (62), leaving just $\mathbbm{1}\otimes S(v_{0},t)$. That is, we define a continuous function
$E:\Omega\mathrm{QCA}_{d}\times[0,1]$ $\to\Omega\mathrm{QCA}_{d}$ (63)
$(\gamma(t),s)$ $\mapsto E(\gamma,t,s)$ (64)
such that $E(\gamma,t,0)=\gamma(t)\otimes\mathbbm{1}$ and $E(\gamma,t,1)=\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$. The deformations all take the form of swapping the positions of gates in the circuit representation of Eq. (62) (Fig. 4). The $S(v_{0}^{\gamma},t)$ factor is a spectator, and so we will focus on $\gamma(t)\otimes S(v_{0},t)^{\dagger}$.
First, in $s\in[0,\tfrac{1}{2}]$, we swap the gates $u_{x^{\prime},x^{\prime}+1}^{\dagger}(v^{\gamma}_{0},t)$ (we denote ancillae sites with a prime, $x^{\prime}$) from the second tensor factor—the second layer—to the first, such that the new gates in the first layer become
$w^{\prime}_{x,x+1}(t)=u^{\dagger}_{x,x+1}(v^{\gamma}_{0},t)w_{x,x+1}(t).$ (65)
This also requires reversing the order of multiplication in
$S^{\dagger}=\left(\prod_{x^{\prime}\in 2{\mathbb{Z}}}u^{\dagger}_{x^{\prime},x^{\prime}+1}\right)\left(\prod_{x^{\prime}\in 2{\mathbb{Z}}+1}u^{\dagger}_{x^{\prime},x^{\prime}+1}\right),$ (66)
so that it matches the circuit structure of $\gamma(t)$. This can be done by adding the gates $\mathrm{SWAP}_{x,x^{\prime}}(2s)$ to the positions marked in Fig. 4(a).
Observe that
$w^{\prime}_{0,1}(1)=\mathbbm{1}\otimes v^{\gamma}_{0}v^{\gamma\dagger}_{1}\text{ and }w^{\prime}_{-1,0}(1)=v^{\gamma\dagger}_{0}v^{\gamma}_{-1}\otimes\mathbbm{1}$ (67)
both act as the identity on site $x=0$ at $t=1$. Thus, in $s\in[\tfrac{1}{2},\tfrac{3}{4}]$, we can introduce $\mathrm{SWAP}(4(s-\tfrac{1}{2}))$ gates to the positions shown in Fig. 4(b) which “roll up” the circuit—the gates previously acting on site 00 now act on-site at $x=\pm 1$ between ancillae. Explicitly, for $s\in[\tfrac{1}{2},\tfrac{3}{4})$, the gate in $E(\gamma,t,s)$ acting across the $(00^{\prime},11^{\prime})$ edge is
$\mathrm{SWAP}_{0,1^{\prime}}(4(s-\tfrac{1}{2}))w^{\prime}_{0,1}(t)\mathrm{SWAP}_{0,1^{\prime}}(4(s-\tfrac{1}{2}))^{\dagger}.$ (68)
As $w^{\prime}_{0,1}(1)$ acts as the identity on site 0, the introduced swap gates cancel at $t=1$, and the entire circuit remains a loop for all $s$. By $s=\tfrac{3}{4}$, the gate acts as the identity on site 0 for all $t\in[0,1]$. We repeat an analogous deformation on the other gate with support on site 00. The part of the circuit which is supported on $x=0$ is thus rolled up to act on the ancillae at sites $x=\pm 1$.
This rolling up procedure can be iterated. At $s=\frac{3}{4}$, we see that the gate $w^{\prime}_{1,2}(t)w^{\prime}_{1^{\prime},1}(t)$ is the identity on sites $1$ and $1^{\prime}$ when $t=1$. Thus, by adding ancillae to site $2$, this gate can also be rolled up within $s\in[\tfrac{3}{4},\tfrac{7}{8}]$ [Fig. 4(c)]. We continue rolling up sites $x$ and $-x$ within $s\in[1-2^{-x-1},1-2^{-x-2}]$, such that, as $s\to 1$, we have $E(\gamma,t,s)\to\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ on any finite patch of the lattice. Technically, this makes $E(\gamma,t,s)$ continuous in the strong operator topology.
The deformation is completed by adding a parallel layer of swap gates which takes $\mathbbm{1}\otimes S(v_{0}^{\gamma},t)$ to $S(v_{0}^{\gamma},t)\otimes\mathbbm{1}\in\mathrm{im}(S)$. If the circuit decomposition Eq. (61) gives $w_{x,x+1}=u_{x,x+1}(v)$ [as it does when using the $\mathrm{Tr}_{(x,\infty)\times{\mathbb{Z}}^{d}}{H(t)}$ construction of Appendix B.1] then $E$ is a deformation retract, and Eq. (51) holds.
Appendix C Homotopy groups of QCA
The $\Omega$-spectrum property reduces the calculation of homotopy groups $\pi_{n}(\mathrm{QCA}_{d})$ of QCA to either finding the connected components $\pi_{0}(\mathrm{QCA}_{d-n})$ or the homotopy groups of a zero-dimensional QCA, $\pi_{n-d}(\mathrm{QCA}_{0})$. In this appendix, we summarize progress in the literature in computing $\pi_{0}(\mathrm{QCA}_{d-n})$ for spin systems (Appendix C.1), and compute all the groups $\pi_{n-d}(\mathrm{QCA}_{0})$ (Appendix C.2). We briefly comment on the classifcation of QCA on lattices other than ${\mathbb{Z}}^{d}$ in Appendix C.3, but leave any technical arguments to Ref. [107].
Unlike the previous two appendices, the results of this appendix are specific for spin systems. The homotopy groups of the space of fermionic QCA will differ from the groups presented here.
C.1 Connected components
In general, the connected components $\pi_{0}(\mathrm{QCA}_{d})$ are not known. However, these groups have been calculated for $d\in\{1,2\}$ (and $d=0$, which is trivial), and the groups are known to be nontrivial in some higher dimensions.
For one-dimensional QCA, the complete classification [that is, $\pi_{0}(\mathrm{QCA}_{1})$] was obtained by Ref. [58], which also explicitly defined an invariant which diagnoses which connected component a given QCA belongs to. All one-dimensional QCA are path connected to a shift—a QCA which acts as a translation on some subset of the degrees of freedom. For example, in a chain of qubits, each QCA is connected to a QCA which translates all qubits by $r$ sites either right or left. For qubit chains, QCA are classified by the group ${\mathbb{Z}}$.
However, we stabilize QCA by allowing ancillae of any Hilbert space dimension to be added to any site by tensor product. Thus, we must assume that the chain also has qutrit (three-state) degrees of freedom, and other higher level qudits. While a 4-level qudit can be decomposed into two qubits, and so does not give new invariants compared to the qubit system, qutrits and other prime qudits all produce invariants. A shift of qutrits cannot be connected to a shift of qubits.
The resulting classifying group is succinctly presented as
$\pi_{0}(\mathrm{QCA}_{1})\cong\log{\mathbb{Q}}^{\times},$ (69)
where ${\mathbb{Q}}^{\times}$ is the set of positive rationals with multiplication as the group operation, and the group operation in $\log{\mathbb{Q}}^{\times}$ is addition. Connecting to our discussion, this can be written as
$\log{\mathbb{Q}}^{\times}\cong\bigoplus_{p\text{ prime}}{\mathbb{Z}}\log p.$ (70)
Each prime Hilbert space dimension is labelled by $\log p$, and the integer coefficients give the shift index for that qudit species.
The classification of two-dimensional QCA on spheres was shown to be trivial in Ref. [61],
$\pi_{0}(\mathrm{QCA}_{2})\cong 0.$ (71)
There is some subtlety in this statement. For QCA on a general two-dimensional manifold—that is, the lattice is obtained from a triangulation of a manifold—the classification may be nontrivial. However, all invariants relate to one-dimensional cycles in the manifold, and so belong to a lower dimensional part of the classification. For appropriate choices of topology on $\mathrm{QCA}_{2}$ [107] (where the manifold is ${\mathbb{R}}^{2}$), this allows all two-dimensional QCA on the plane to be connected by continuous paths.
The classification of three dimensional QCA is still open. It is conjectured that $\pi_{0}(\mathrm{QCA}_{3})$ is isomorphic to the Witt group of modular tensor categories [56, 108, 55, 109]—that is, it has the same group structure as stacking two-dimensional topological orders, modulo an equivalence relation which identifies phases which may share a gapped edge. Several nontrivial components of $\pi_{0}(\mathrm{QCA}_{3})$ have been identified, which are believed to form a subgroup [55],
$\pi_{0}(\mathrm{QCA}_{3})\geq{\mathbb{Z}}_{8}\oplus{\mathbb{Z}}_{2}.$ (72)
The elements of this subgroup can, indeed, be identified with two-dimensional topologically ordered phases. In fact, the ALT phases corresponding to these elements exhibit that topological order (specifically, the associated anyonic excitations) on their two dimensional surfaces [48, 56, 55, 109].
C.2 Zero dimensions
The space of zero dimensional QCA with finite Hilbert space dimension, say $N$, is naturally identified with the projective unitary group,
$\mathrm{QCA}_{0}(N)\cong PU(N).$ (73)
The complication in finding the homotopy groups of $\mathrm{QCA}_{0}$ is the stabilization by adding ancillae. That is, the $N\to\infty$ limit.
C.2.1 Unstable invariants
For finite but large $N$, the homotopy groups of $PU(N)$ are all known,
$\pi_{n}(PU(N))\cong\left\{\begin{array}[]{l l}{\mathbb{Z}}_{N}&\quad\text{for }n=1,\\ 0&\quad\text{for }n<2N\text{ and even,}\\ {\mathbb{Z}}&\quad\text{for }n<2N\text{ and odd.}\end{array}\right.$ (74)
There are even integral formulae for winding number invariants indexing the elements of these groups. Given any continuous map $V:S^{n}\to PU(N)$ (where $S^{n}$ is the $n$-dimensional sphere) it is possible to find a lift $\tilde{V}:S^{n}\setminus\mathrm{pt}\to U(N)$ from the punctured sphere ($\mathrm{pt}$ is the topological space with a single point) to the usual unitary group. That is, if $q:U(N)\to PU(N)$ is the quotient map from the unitary group to the projective unitaries, we have $q\tilde{V}=V|_{S^{n}\setminus\mathrm{pt}}$. This is possible because the punctured sphere is contractible, and so it is possible to give a smooth gauge to $\tilde{V}$ away from the puncture. Defining $A=(\mathrm{d}\tilde{V})\tilde{V}^{\dagger}$, where $\mathrm{d}$ is the exterior derivative on $S^{n}\setminus\mathrm{pt}$, the winding number invariants are given by [31, 130]
$W_{n}[\tilde{V}]=C_{n}\int_{S^{n}\setminus\mathrm{pt}}\mathrm{Tr}\left[A^{\wedge n}\right],$ (75)
where $C_{n}$ is a constant independent of $N$ and $A^{\wedge n}=A\wedge\cdots\wedge A$ is the $n$-fold wedge product of $A$ with itself. For even $n$ we have $W_{n}=0$, while for odd $n\geq 3$ the winding number is an integer which only depends on the homotopy class of $V$, so we can denote them $W_{n}[V]$. For $n=1$, only
$w_{1}[V]=W_{1}[\tilde{V}]\bmod N\in{\mathbb{Z}}_{N}$ (76)
acts as an invariant for $V$. Indeed, redefining $\tilde{V}\to e^{i\theta}\tilde{V}$ does not affect $V$, but takes $W_{1}\mapsto W_{1}+N$.
C.2.2 Stable invariants
The complication is that these winding numbers are invariant under the direct sum with an ancilla, not the tensor product. Indeed, denoting $\mathbbm{1}_{M}$ as the identity in $PU(M)$, we have from Eq. (75)
$W_{n}[V\otimes\mathbbm{1}_{M}]=MW_{n}[V].$ (77)
More generally, for $Q\in PU(M)$,
$W_{n}[V\otimes Q]=MW_{n}[V]+NW_{n}[Q].$ (78)
To find invariants which are stable under the addition of ancillae, we simply need to rescale the winding numbers as
$\tilde{W}_{n}[V]=C_{n}\int_{S^{n}\setminus\mathrm{pt}}\frac{\mathrm{Tr}\left[A^{\wedge n}\right]}{\mathrm{Tr}\left[\mathbbm{1}\right]}.$ (79)
These are no longer integer invariants, but rather rational invariants [modulo 1 for $n=1$]. Then Eq. (78) shows that $\tilde{W}_{n}[V\otimes Q]=\tilde{W}_{n}[V]+\tilde{W}_{n}[Q]$.
C.2.3 Classification
At a more formal level, for all $N,M\in{\mathbb{N}}$ we have injective homomorphisms
$\iota_{N\to NM}:PU(N)$ $\to PU(NM)$ (80)
$V$ $\mapsto V\otimes\mathbbm{1}_{M}$
which define a directed system of topological groups. $\mathrm{QCA}_{0}$ is the direct limit
[…]
with respect to this system. Strictly, we require that the multiplicity $n_{p}$ of every prime $p$ in the factorization $N=2^{n_{2}}3^{n_{3}}\ldots p^{n_{p}}\ldots$ approaches infinity, $n_{p}\to\infty$. The homotopy groups of $\mathrm{QCA}_{0}$ can then be computed from the induced system
$\iota^{*}_{N\to NM}:\pi_{n}(PU(N))$ $\to\pi_{n}(PU(NM))$ (82)
$W_{n}[V]$ $\mapsto MW_{n}[V].$
By rescaling the winding numbers, we find that
[…]
$\cong\left\{\begin{array}[]{l l}{\mathbb{Q}}/{\mathbb{Z}}&\quad\text{for }n=1,\\ 0&\quad\text{for }n\text{ even,}\\ {\mathbb{Q}}&\quad\text{for }n\text{ odd.}\end{array}\right.$ (87)
The group operation is given by the usual addition of rational numbers, and corresponds to the tensor product of QCA.
C.2.4 Limited ancillae
We obtained a classification by rationals because we allowed ancillae of any dimension to be appended to the system. Thus, all Hilbert space dimensions can occur in the denominator of $\tilde{W}_{n}$. If, instead, we considered only a limited collection of Hilbert space dimensions for the ancillae, we would obtain a different classification. In fact, this classification will be a useful warm-up for the case of QCA with symmetry.
Say our available ancillae dimensions are given by
=$\mathcal{R}=\{\rho\in{\mathbb{N}}\,:\,\text{ancilla of dimension }\rho\text{ exists}\},$ (88)
and the system is built out of degrees of freedom with these dimensions. Then $\mathcal{R}$ forms a submonoid (similar to a group, except without inverses) of ${\mathbb{N}}$ under multiplication, as if $\mathbbm{1}_{\rho}$ and $\mathbbm{1}_{\sigma}$ can be ancillae, then so can $\mathbbm{1}_{\rho\sigma}=\mathbbm{1}_{\rho}\otimes\mathbbm{1}_{\sigma}$. $\mathcal{R}$ also contains $1$, which corresponds to adding no ancillae.
A similar directed system to Eqs. (8081) defines the limit $\mathrm{QCA}_{0}^{\mathcal{R}}$ describing the stabilization of QCA by the ancillae in $\mathcal{R}$. The only difference is that now the system proceeds through increasing $\rho\in\mathcal{R}$, rather than $N\in{\mathbb{N}}$. Thus, not all denominators may occur in the rescaled winding numbers. Defining the algebraic localization (not to be confused with localization in physical space)
$\mathcal{R}^{-1}{\mathbb{Z}}:=\left\{\frac{p}{q}\in{\mathbb{Q}}\,:\,p\in{\mathbb{Z}},q\in\mathcal{R}\right\},$ (89)
then $\mathcal{R}^{-1}{\mathbb{Z}}$ is an additive group (indeed, a ring) and we have
$\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{0})\cong\left\{\begin{array}[]{l l}(\mathcal{R}^{-1}{\mathbb{Z}})/{\mathbb{Z}}&\quad\text{for }n=1,\\ 0&\quad\text{for }n\text{ even,}\\ \mathcal{R}^{-1}{\mathbb{Z}}&\quad\text{for }n\text{ odd.}\end{array}\right.$ (90)
C.3 QCA on manifolds
The $\Omega$-spectrum property discussed in Appendix B assumes that QCA are defined on the integer lattice ${\mathbb{Z}}^{d}$. By coarse graining, any locally finite set of points in ${\mathbb{R}}^{d}$ can be deformed to ${\mathbb{Z}}^{d}$, so the QCA is more properly based on the manifold ${\mathbb{R}}^{d}$. This distinction is important, as the classification of QCA based on the $d$-dimensional manifold $M$, $\mathrm{QCA}_{M}$, depends on the homology of $M$ [61] and not just its dimension.
It is natural to conjecture that QCA on an orientable manifold $M$ can all be patched together from QCA defined on the open charts in an atlas for $M$. That is, we construct a map
$\mathrm{QCA}_{M}\to\left[M,\mathrm{QCA}_{d}\right]$ (91)
(where $\left[M,\mathrm{QCA}_{d}\right]$ is the space of continuous functions from $M$ to $\mathrm{QCA}_{d}$) by identifying the action of $V\in\mathrm{QCA}_{M}$ near each point $x\in M$ as a QCA on ${\mathbb{R}}^{d}$ via a chart in $M$ containing $x$. We suspect that a homotopy equivalence $\mathrm{QCA}_{M}\simeq\left[M,\mathrm{QCA}_{d}\right]$ can be constructed in this way.
This leads to the hypothesis that, for orientable $M$ [116],
$\pi_{0}\left[X,\mathrm{QCA}_{M}\right]\cong\pi_{0}\left[X\times M,\mathrm{QCA}_{d}\right]\cong h^{d}(X\times M).$ (92)
We defer the majority of work to Ref. [107], but we remark here that Eq. (92) reproduces the classification of QCA on two-dimensional orientable manifolds in Ref. [61]. For non-orientable manifolds, one must consider twisted generalized cohomology—that is, $\mathrm{QCA}_{M}$ would be homotopy equivalent to spaces of sections in a $\mathrm{QCA}_{d}$-bundle over $M$ [103, 104].
Appendix D Homotopy groups of symmetric QCA
The QCA formalism for classifying localized phases can be adapted to incorporate symmetry constraints, and thus provide a classification of symmetry-enriched ALT phases (SALT phases), and more generally of symmetry-enriched MBL phases with short-ranged entanglement. In this appendix, we show that the $\Omega$-spectrum property of QCA respects any on-site unitary group action, from which it follows that symmetric QCA also form an $\Omega$-spectrum (Appendix D.1). While the $\Omega$-spectrum property of invertible states is linked to a classification of symmetry-enriched phases by $BG$-parameterized families of states ($BG$ is the classifying space for the symmetry group), this construction fails for QCA (Appendix D.2). Nonetheless, we identify complete invariants for all homotopy groups of symmetric QCA for $d\leq 1$, and give a presentation of the group structure (Appendices D.3D.4).
An important observation is that the classification of symmetric QCA depends on the representation of the symmetry action [93, 26, 114]. This can be formalized within the stabilization procedure of QCA. Each added ancilla carries some representation of the symmetry group $G$. We do not assume that all representations of $G$ are necessarily available. Define a countable set of representations which occur in the ancillae,
:=$\mathcal{R}:=\{\rho\in\mathrm{Rep}(G)\,:\,\text{some ancillae carry }\rho\},$ (93)
where $\mathrm{Rep}(G)$ is the tensor category of finite dimensional unitary representations of $G$. The set $\mathcal{R}$ is a multiplicative submonoid of $\mathrm{Rep}(G)$ under tensor product, as if $\rho$ and $\sigma$ are representations carried by some ancillae, then $\rho\otimes\sigma$ is the representation of the ancilla formed by stacking both ancillae. $\mathcal{R}$ also contains the trivial one-dimensional representation $1$, which corresponds to appending no ancillae. This definition of $\mathcal{R}$ generalizes Eq. (88), for which we can regard each ancilla as a representation of the trivial group. (One can compare $\mathcal{R}$ to the closely related notion of a $G$-universe [131], which is defined similarly, except that a $G$-universe is usually required to be closed under direct sum, rather than tensor product.)
We say that two QCA $V_{1}$ and $V_{2}$ which carry a group action by $G$ are $\mathcal{R}$-equivalent if ancillae from $\mathcal{R}$ can be appended to each of them such that they become equal, and carry the same representation. In the literature, two QCA have been called strongly equivalent if they are connected by a symmetric path modulo $\mathcal{R}$-equivalence [93, 114].
D.1 Symmetric swap
In Appendix B, we argued that the swindle map $S:\mathrm{QCA}_{d-1}\to\Omega\mathrm{QCA}_{d}$ is a homotopy equivalence. Both the swindle map and the deformations which demonstrated this equivalence were defined in terms of a path $\mathrm{swap}_{jk}(t)$ from the identity map to a swap gate between two sites with the same local Hilbert space dimension. In the context of QCA with an on-site $G$-action, we can repeat all the same arguments, provided we can find a path $\mathrm{swap}^{\rho}_{jk}(t)$ which commutes with the representation $\rho$.
Consider two sites with Hilbert space $\mathcal{H}_{j}\otimes\mathcal{H}_{k}$ and a representation $\rho\otimes\rho$ of the group $G$. It is a well known fact that the representation $\rho\otimes\rho$ decomposes into a symmetric and alternating subspace. Indeed, we have
$\mathrm{swap}_{jk}(1)\rho_{g}\otimes\rho_{g}=\rho_{g}\otimes\rho_{g}\mathrm{swap}_{jk}(1)$ (94)
for all $g\in G$. Thus, all $\rho_{g}\otimes\rho_{g}$ become block diagonal in an eigenbasis of the swap gate, $\mathrm{swap}_{jk}(1)$. Decomposing
$\mathcal{H}_{j}\otimes\mathcal{H}_{k}\cong\mathrm{Sym}(\mathcal{H}_{j}\otimes\mathcal{H}_{k})\oplus\mathrm{Alt}(\mathcal{H}_{j}\otimes\mathcal{H}_{k}),$ (95)
where $\mathrm{Sym}(\mathcal{H}_{j}\otimes\mathcal{H}_{k})$ is the $(+1)$-eigenspace of $\mathrm{swap}_{jk}(1)$ and $\mathrm{Sym}(\mathcal{H}_{j}\otimes\mathcal{H}_{k})$ its $(-1)$-eigenspace, we have
$\rho\otimes\rho\cong\rho_{\mathrm{Sym}}\oplus\rho_{\mathrm{Alt}}.$ (96)
Then the path of operators
$\mathrm{swap}^{\rho}_{jk}(t)=\mathbbm{1}_{\mathrm{Sym}}\oplus(e^{i\pi t}\mathbbm{1}_{\mathrm{Alt}})$ (97)
is continuous, connects $\mathbbm{1}$ to the swap gate, and commutes with $\rho_{g}\otimes\rho_{g}$ for all $g\in G$ and $t\in[0,1]$.
Then, given some on-site group action $\rho$ on $\mathrm{QCA}_{d-1}\ni v$, we can construct a loop of QCA on a doubled Hilbert space
$u^{\rho}_{x,x+1}(v,t):=v_{x}\mathrm{SWAP}^{\rho}_{x,x+1}(t)v_{x}^{\dagger}\mathrm{SWAP}^{\rho}_{x,x+1}(t)^{\dagger},$ (98)
where $\mathrm{SWAP}^{\rho}_{x,x+1}(t)$ is a product of $\mathrm{swap}^{\rho}_{jk}(t)$ paths, as in Eq. (6). Multiplying these together gives an instance of the swindle map
$S^{\rho}(v,t)=\left(\prod_{x\in 2{\mathbb{Z}}+1}u^{\rho}_{x,x+1}(v,t)\right)\left(\prod_{x\in 2{\mathbb{Z}}}u^{\rho}_{x,x+1}(v,t)\right),$ (99)
which respects the group action in the sense that
$\left(\prod_{x}\rho_{x}\right)S^{\rho}(v,t)\left(\prod_{x}\rho^{\dagger}_{x}\right)=S^{\rho}(\rho v\rho^{\dagger},t).$ (100)
That is, the group action on the $d$-dimensional loop agrees with the action on the $(d-1)$-dimensional QCA. Further, as the deformation from $\Omega\mathrm{QCA}_{d}$ to the image of $S(v,t)$ in Appendix B proceeded exclusively through interleaving finite depth circuits with swap gates, we can make all the same deformations with $\mathrm{SWAP}^{\rho}_{x,x+1}(t)$. Thus, $S^{\rho}$ gives a $G$-homotopy equivalence—a homotopy equivalence which respects the $G$ action.
This makes $\mathrm{QCA}$ a $G$-spectrum, or more specifically a (naive) $\Omega$-spectrum with $G$-action. Formally, this defines a (Bredon) equivariant cohomology theory. On general grounds, it is known that the fixed point subspaces of the spectrum—QCA which commute with the group action—themselves form an $\Omega$-spectrum [131]. (An equivariant map must map fixed points to fixed points.) We write the fixed point spaces as $\mathrm{QCA}_{d}^{\mathcal{R}}$, where we use $\mathcal{R}$ as the decoration rather than $\rho$ to denote the stabilization by appending ancillae. Also define $\mathrm{QCA}_{d}^{\mathcal{R}}$ for negative $d$ by $\mathrm{QCA}_{d}^{\mathcal{R}}=\Omega^{-d}\mathrm{QCA}_{0}^{\mathcal{R}}$. Then we have
$\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})\cong\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{d-n}).$ (101)
More generally, there is a generalized cohomology theory
$h^{d}_{\mathcal{R}}(X):=\pi_{0}\left[X,\mathrm{QCA}^{\mathcal{R}}_{d}\right]$ (102)
which classifies parameterized families of symmetric QCA.
D.2 Symmetric QCA are not classified by $BG$-parameterized QCA
Invertible states are also conjectured to form an $\Omega$-spectrum [96, 95, 102, 103, 104]. Further, it is believed that the classification of $G$-symmetry-enriched invertible states can be deduced just from the cohomology theory $h^{d}_{\mathrm{inv}}(X):=\pi_{0}\left[X,\mathrm{inv}_{d}\right]$ and the abstract group $G$. In particular, denoting $\mathrm{inv}^{G}_{d}$ for the space of fixed points under the $G$-action, it is conjectured that
$\pi_{0}(\mathrm{inv}^{G}_{d})\cong h^{d}_{\mathrm{inv}}(BG),$ (103)
where $BG$ is the classifying space for the group $G$. That is, connected components of symmetric QCA are in bijection with families of not-necessarily-symmetric QCA parameterized by $BG$. This conjecture can has a natural interpretation in which the map in $\pi_{0}\left[BG,\mathrm{inv}_{d}\right]$ encodes the fusion rules of domain walls [96, 95, 113, 102, 103, 104].
The corresponding hypothesis for QCA is, in general, false:
$\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{d})\not\cong h^{d}(BG).$ (104)
Indeed, the hypothetical classification by $h^{d}(BG)$ is independent of the symmetry representation $\rho$, and as the strong classification of QCA (with restricted representations of ancillae) is known to depend on the representation, it cannot be the case that symmetric QCA are classified by $h^{d}(BG)$. In fact, Ref. [26] computed the classification of $U(1)$-symmetric QCA in one dimension with an arbitrary symmetry representation. In this case, while the classification depends on the symmetric QCA $V$ and the representation $\rho$, the invariant associated to the pair $(V,\rho)$ does not depend on which stabilizing ancillae can be included, $\mathcal{R}$. Thus, some of the subtleties associated to stacking ancillae are absent in this case. Regardless, we already see the $\rho$-independent $h^{1}(BG)$ cannot be a complete classification.
By explicit computation, we show
$\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{1})\not\geq h^{1}(BU(1))$ (105)
for any collection $\mathcal{R}$ of representations of $U(1)$. Indeed, the left hand side was found in Ref. [26]. We will repeat this classification with greater generality in Appendix D.4, but for now all we need is that the classifying group is free abelian, so no element has finite order.
On the right hand side of Eq. (105), we can compute $h^{1}(BU(1))$ by solving the Atiyah-Hirzebruch spectral sequence (AHSS) [103, 104]. The AHSS is a method for evaluating generalized cohomology theories $h^{d}(X)$ given knowledge of $h^{p}(\mathrm{pt})$ for all $p\leq d+1$ and the cohomology groups of $X$, $H^{p}(X;M)$. In general, it gives that $h^{d}(X)$ is a quotient of a subgroup of
$\bigoplus_{p=0}^{\infty}H^{p}(X;h^{d-p}(\mathrm{pt})).$ (106)
Identifying the right subgroup and quotient is difficult in general, but for $X=BU(1)\simeq{\mathbb{C}}\mathbb{P}^{\infty}$ and $d=1$ it happens to be easy—the spectral sequence stabilizes on the $E_{2}$ page, so the subgroup is the entire direct sum and the quotient is trivial. We have
$h^{1}(BU(1))$ $\cong\bigoplus_{p=0}^{\infty}H^{p}({\mathbb{C}}\mathbb{P}^{\infty};h^{1-p}(\mathrm{pt}))$ (107)
$\cong H^{0}({\mathbb{C}}\mathbb{P}^{\infty};\log{\mathbb{Q}}_{+})\oplus 0\oplus H^{2}({\mathbb{C}}\mathbb{P}^{\infty};{\mathbb{Q}}/{\mathbb{Z}})$
$\qquad\qquad\oplus 0\oplus H^{4}({\mathbb{C}}\mathbb{P}^{\infty};{\mathbb{Q}})\oplus 0\cdots$ (108)
$\cong\log{\mathbb{Q}}^{\times}\oplus{\mathbb{Q}}/{\mathbb{Z}}\oplus{\mathbb{Q}}\oplus{\mathbb{Q}}\oplus\cdots.$ (109)
We used that $BU(1)={\mathbb{C}}\mathbb{P}^{\infty}$ and $H^{n}({\mathbb{C}}\mathbb{P}^{\infty};M)=M$ for even $n$ and is trivial otherwise. In particular, we have $h^{1}(BU(1))\geq{\mathbb{Q}}/{\mathbb{Z}}$, which is not a free group—all its elements have finite order. Thus, $\pi_{0}(\mathrm{QCA}^{\mathcal{R}}_{1})$ cannot contain $h^{1}(BU(1))$.
In the case of invertible states, the classification by $h^{d}_{\mathrm{inv}}(BG)$ is justified through a decorated domain wall construction [103, 104]. We speculate that this fails for QCA because, while QCA form a spectrum, that spectrum is non-connective—it has nontrivial homotopy groups $h^{d}(\mathrm{pt})$ in arbitrarily negative degree $d<0$. Thus, the putative domain wall structure encoded in $h^{d}(BG)$ includes algebraic fusion rules for domain walls which exist in negative dimension, which are not physical. This may also be why single-particle symmetric phases are not classified by $h^{d}_{K}(BG)$ even though the phases without symmetry are classified by the $K$-theory spectrum $h^{d}_{K}$ (which has nontrivial homotopy groups in all even degrees) [100].
It seems plausible that all symmetric QCA can be classified through a decorated domain wall construction. Certainly, all examples in the literature have this structure [91, 26, 93]. However, when correctly implemented, this approach must treat the zero-dimensional defects which terminate such constructions [113] with greater care than $h^{d}(BG)$.
D.3 Zero dimensions with symmetry
As in the case without symmetry, the homotopy groups $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ with $d\leq n$ are determined by $\pi_{n-d}(\mathrm{QCA}^{\mathcal{R}}_{0})$. Thus, we make significant progress on characterizing symmetric QCA in higher dimensions through an in-depth analysis of zero-dimensional QCA with symmetry. Our strategy is the same as the case without symmetry: we identify unstable invariants (at fixed Hilbert space dimension) and subsequently find a stable presentation of these invariants which does not change upon adding ancillae. Finally, we characterize the abstract classifying group.
D.3.1 Unstable invariants
We consider a Hilbert space $\mathcal{H}$ with finite dimension $N$ carrying a continuous unitary representation $\rho:G\to U(N)$ of an arbitrary topological group $G$. Then $\rho$ is a direct sum of irreducible representations (irreps) $r$ each with multiplicity $n_{r}$,
$\rho=\bigoplus_{r\text{ irrep of }G}r\otimes\mathbbm{1}_{n_{r}},$ (110)
and we have
$N=\sum_{r\text{ irrep of }G}n_{r}d_{r},$ (111)
where $d_{r}$ is the dimension of the irrep $r$. In particular, all but finitely many $n_{r}$ are zero.
We first describe the classification of symmetric unitaries, before explaining how this classification is modified for projective unitaries. Any family of unitaries $\tilde{V}:S^{n}\to U(N)$ which commutes with $\rho$ must decompose as
$\tilde{V}=\bigoplus_{r\text{ irrep of }G}\mathbbm{1}_{d_{r}}\otimes\tilde{V}_{r}.$ (112)
Any deformation of this family which continues to commute with $\rho$ must maintain this structure. Thus, the families $\tilde{V}_{1}({\vec{\theta}})$ and $\tilde{V}_{2}({\vec{\theta}})$ are homotopic through a symmetric homotopy if and only if $\tilde{V}_{1r}({\vec{\theta}})$ and $\tilde{V}_{2r}({\vec{\theta}})$ are homotopic for all $r$. We have a complete set of invariants given by the winding numbers of each $\tilde{V}_{r}$,
$W^{\rho}_{n}[\tilde{V}]=(W_{n}[\tilde{V}_{r}])_{r\text{ irrep of }G}.$ (113)
With some foresight, we choose to identify $W^{\rho}_{n}[\tilde{V}]$ with an element of the module
$\mathrm{Rep}(G)[\pi_{n}(U(N))]=\Bigg{\{}\sum_{\rho\in\mathrm{Rep}(G)}\rho\cdot W_{n}^{\rho}\\ :\,W_{n}^{\rho}\in\pi_{n}(U(N))\Bigg{\}},$ (114)
which consists of finite formal sums of winding numbers with coefficients in $\mathrm{Rep}(G)$, regarded as a rig (a ring without negatives). That is, addition in $\mathrm{Rep}(G)$ is direct sum, multiplication is tensor product (which distributes over direct sums) and elements are regarded up to unitary equivalence of representations, so that addition and multiplication are commutative. We do not append the formal negatives used to define the representation ring $R_{G}$. However, as $\pi_{n}(U(N))\in\{0,{\mathbb{Z}}\}$ for sufficiently large $N$, the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is isomorphic to either 00 or $R_{G}$.
Then we define the complete invariant for symmetric unitaries
$W^{\rho}_{n}[\tilde{V}]=\sum_{r\text{ irrep of }G}r\cdot W_{n}[\tilde{V}_{r}].$ (115)
QCA are properly projective unitaries, rather than unitaries. $W^{\rho}_{n}[\tilde{V}]$ continues to function as a complete invariant for $n\geq 2$, as in the case without symmetry. We may thus denote them $W^{\rho}_{n}[V]$, where $V:S^{n}\to PU(N)$ is a family of symmetric projective unitaries which lie in the connected component of the identity. However, there are symmetric projective unitaries $V$ which cannot be expressed as symmetric proper unitaries (requiring a larger group of $n=0$ invariants), and there are loops $V(\theta)$ of symmetric proper unitaries which become equivalent when regarded as symmetric projective unitaries (requiring a smaller group of $n=1$ invariants).
We first recall that connected components of symmetric projective unitaries [$n=0$] are labelled by one-dimensional representations of $G$. Given a symmetric projective unitary $V$, we can arbitrarily choose a proper unitary $\tilde{V}$ in the preimage of $V$ under the quotient by global phases. The fact that $V$ is symmetric only implies that $\tilde{V}$ is symmetric up to a global phase,
$\rho(g)^{\dagger}\tilde{V}\rho(g)=\chi[V](g)\tilde{V}.$ (116)
Note that $\chi[V]$ is the same when $\tilde{V}$ is multiplied by a global phase, justifying the notation $\chi[V]$ rather than $\chi[\tilde{V}]$. Using that $\rho$ is a representation, we see that $\chi[V]:G\to U(1)$ is a one-dimensional representation of $G$. From Eq. (110), we see that $\tilde{V}$ must decompose as
$\tilde{V}=\sum_{r\text{ irrep of }G}\sigma_{r\to\chi[V]r}\otimes\tilde{V}_{r},$ (117)
where $\sigma_{r\to\chi[V]r}$ is a matrix with a single nonzero element equal to 1, such that $\tilde{V}$ maps the irrep $r$ block of $\rho$ to the irrep $\chi[V]r$ block. Any path of projective unitaries $V(t)$ lifts to a path of proper unitaries $\tilde{V}(t)$ (as the unit interval is contractible), and the space of permutations $\sigma_{r\to\chi[V]r}$ is discrete, so the permutation label $\chi[V]$ is an invariant for path components of projective unitaries. Any unitaries with the same $\chi[V]$ in the decomposition Eq. (117) can be connected by a path, simply by connecting each $\tilde{V}_{r}$ by a path. Thus, $\chi[V]$ is a complete invariant.
However, not all $\chi[V]\in\mathrm{Hom}(G\to U(1))=:X_{G}$ can occur in a finite Hilbert space. For the $\tilde{V}$ in Eq. (117) to be unitary, we need that $\sum_{r}\sigma_{r\to\chi[V]r}$ is a finite-dimensional permutation matrix, and so in particular $\chi[V]$ has finite order: there is an $n$ such that $\chi[V](g)^{n}=1$ for all $g$. Thus, $\chi[V]$ must belong to the torsion subgroup of $X_{G}$ (the subgroup of elements of finite order, with the group operation in $X_{G}$ defined pointwise), which we denote $X_{G}^{T}$. Conversely, given any $\chi[V]\in X_{G}^{T}$ with order $n$, we can take
$\rho=\bigoplus_{p=0}^{n-1}\chi[V]^{p}$ (118)
and then the unitary $\sum_{p=0}^{n-1}\sigma_{\chi[V]^{p}\to\chi[V]^{p+1}}$ achieves the invariant $\chi[V]$.
As an aside related to the discussion in Appendix D.2, we note that
$X_{G}^{T}\cong H^{1}(G;{\mathbb{Q}}/{\mathbb{Z}})\cong H^{1}(G;\pi_{1}(\mathrm{QCA}_{0}))$ (119)
for the vast majority of groups $G$ relevant to physics. For instance, a sufficient condition is that the component group $G/G^{0}$ (where $G^{0}$ is the identity component of the topological group $G$) is finitely generated.
For $n=1$, we can multiply a loop of proper unitaries $\tilde{V}(\theta)$ by $e^{im\theta}$ and change all the winding numbers, while still forming the same loop of projective unitaries after the quotient by global phases. Indeed, we find
$W^{\rho}_{1}[e^{im\theta}\tilde{V}]$ $=\sum_{r\text{ irrep of }G}r\cdot W_{1}[e^{im\theta}\tilde{V}_{r}]$ (120)
$=\sum_{r\text{ irrep of }G}r\cdot(mn_{r}+W_{1}[\tilde{V}_{r}])$ (121)
$=\left(\sum_{r\text{ irrep of }G}r\otimes\mathbbm{1}_{n_{r}}\right)\cdot m+W^{\rho}_{1}[\tilde{V}]$ (122)
$=\rho\cdot m+W^{\rho}_{1}[\tilde{V}],$ (123)
where we used
$r\cdot n_{r}=r\cdot(1+\cdots+1)=r+\cdots+r=r\otimes\mathbbm{1}_{n_{r}}\cdot 1$ (124)
for $n_{r}\geq 0$ and the decomposition of $\rho$ into irreps, $\rho=\sum_{r}r\otimes\mathbbm{1}_{n_{r}}$. Thus, we must consider the coset
$w^{\rho}_{1}[V]=W^{\rho}_{1}[\tilde{V}]+\rho\cdot\pi_{1}(U(N))\cong W^{\rho}_{1}[\tilde{V}]+\rho\cdot{\mathbb{Z}},$ (125)
which is now a well-defined invariant for loops of projective unitaries $V(\theta)$.
D.3.2 Stable invariants
With the unstable invariants identified, we must now consider the stabilization process of appending ancillae. We have a directed system among symmetric projective unitary groups $PU(N)^{\rho}$ [the subset of $PU(N)$ which commutes with the representation $\rho$] given by
$\iota_{\rho\to\rho\sigma}:PU(N)^{\rho}$ $\to PU(NM)^{\rho\sigma},$ (126)
$V$ $\mapsto V\otimes\mathbbm{1}$
where we take $\rho,\sigma\in\mathcal{R}$, $N=\mathrm{dim}(\rho)$, and $M=\mathrm{dim}(\sigma)$. The stabilized space $\mathrm{QCA}^{\mathcal{R}}_{0}$ is the direct limit
[…]
Strictly, we require that the representation $\rho$ grows through the directed system such that each representation in $\mathcal{R}$ eventually occurs as a tensor factor in $\rho$.
Starting at $n=0$, we find the invariant $\chi[V]=e^{i\omega^{\rho}[V]}$ of $\pi_{0}(PU(N)^{\rho})$ is actually already stable to appending ancillae. With $\omega^{\rho}[V]\in{\mathbb{R}}/(2\pi{\mathbb{Z}})\cong U(1)$ defined modulo $2\pi$, we can see from Eq. (116) that
$\iota_{\rho\to\rho\sigma}^{*}(\omega^{\rho}[V])=\omega^{\rho\sigma}[V\otimes\mathbbm{1}]=\omega^{\rho}[V]$ (128)
and more generally
$\omega^{\rho\sigma}[V\otimes Q]=\omega^{\rho}[V]+\omega^{\sigma}[Q].$ (129)
To compute the action of the induced maps on the remaining unstable invariants, $\iota^{*}_{\rho\to\rho\sigma}(W_{n}^{\rho}[V])=W_{n}^{\rho\sigma}[V\otimes\mathbbm{1}]$, we suppose that we have Clebsch-Gordan multiplicities $C_{rr^{\prime}}^{k}\in{\mathbb{Z}}$ such that, for $r,r^{\prime},k$ irreducible,
$rr^{\prime}=\sum_{k}C^{rr^{\prime}}_{k}k\in\mathrm{Rep}(G),$ (130)
where $C^{rr^{\prime}}_{k}k=k+\cdots+k=\mathbbm{1}_{C^{rr^{\prime}}_{k}}\otimes k$. Taking
$\rho=\sum_{r}n^{\rho}_{r}r\quad\text{and}\quad\sigma=\sum_{r^{\prime}}n^{\sigma}_{r^{\prime}}r^{\prime},$ (131)
we have, with summation implied,
$\rho\sigma=n_{r}^{\rho}n_{r^{\prime}}^{\sigma}C_{k}^{rr^{\prime}}k.$ (132)
Then $V\otimes\mathbbm{1}$ decomposes into the irreducible subspaces of $\rho\sigma$ (possibly after multiplication by a permutation of the irrep subspaces) as
$V\otimes\mathbbm{1}\cong\bigoplus_{k}\left[\mathbbm{1}_{d_{k}}\otimes\left(\bigoplus_{r,r^{\prime}}V_{r}\otimes\mathbbm{1}_{n_{r^{\prime}}^{\sigma}C^{rr^{\prime}}_{k}}\right)\right],$ (133)
and the invariant becomes
$W^{\rho\sigma}_{n}[V\otimes\mathbbm{1}]$ $=\sum_{k,r,r^{\prime}}k\cdot W_{1}[V_{r}\otimes\mathbbm{1}_{n_{r^{\prime}}^{\sigma}C^{rr^{\prime}}_{k}}]$ (134)
$=\sum_{k,r,r^{\prime}}(n_{r^{\prime}}^{\sigma}C^{rr^{\prime}}_{k}k)\cdot W_{1}[V_{r}]$ (135)
$=\sum_{r}(\sigma r)\cdot W_{1}[V_{r}]$ (136)
$=\sigma\cdot W^{\rho}_{n}[V].$ (137)
An analogous calculation produces the generalization
$W^{\rho\sigma}_{n}[V\otimes Q]=\sigma\cdot W^{\rho}_{n}[V]+\rho\cdot W^{\sigma}_{n}[Q].$ (138)
Then the induced directed system is
$\iota^{*}_{\rho\to\rho\sigma}:\mathrm{Rep}(G)[\pi_{n}(U(N))]$ $\to\mathrm{Rep}(G)[\pi_{n}(U(NM))]$
$W^{\rho}_{n}[V]$ $\mapsto\sigma\cdot W^{\rho}_{n}[V].$ (139)
We recover Eq. (82) by taking $G=0$ to be the trivial group.
Thus, we can identify a stable invariant by formally dividing $W^{\rho}_{n}[V]$ by the representation $\rho$,
$\tilde{W}^{\rho}_{n}[V]=\frac{W^{\rho}_{n}[V]}{\rho}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))],$ (140)
which then obeys $\tilde{W}^{\rho\sigma}_{n}[V\otimes\mathbbm{1}]=\tilde{W}^{\rho}_{n}[V]$. Here, $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$ is the algebraic localization of the module $\mathrm{Rep}(G)[\pi_{n}(U(N))]$, defined in the next section.
For $n=1$, the invariant is a coset
$\tilde{w}^{\rho}_{1}[V]=\tilde{W}^{\rho}_{1}[V]+1\cdot{\mathbb{Z}}\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot{\mathbb{Z}}).$ (141)
D.3.3 Classification
If $M$ is a module over $R$ and $S\subseteq R$ is a multiplicatively closed subset of $R$, then the algebraic localization of $M$ is the set $S^{-1}M:=(M\times S)/\sim$, where the equivalence relation is
$(m_{1},s_{1})\sim(m_{2},s_{2})\iff\exists\sigma\in S\text{ such that}\\ \sigma\cdot(s_{1}\cdot m_{2}-s_{2}\cdot m_{1})=0\in M.$ (142)
We write the equivalence class of $(m,s)$ as $m/s$ and equip this set with the addition operation and $R$-action
$\frac{m_{1}}{s_{1}}+\frac{m_{2}}{s_{2}}$ $=\frac{s_{2}\cdot m_{1}+s_{1}\cdot m_{2}}{s_{1}s_{2}},$ (143)
$\rho\cdot\frac{m}{s}$ $=\frac{\rho\cdot m}{s}.$ (144)
The formal ratios $\tilde{W}^{\rho}_{n}[V]$ are properly defined as elements of the localization $\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{n}(U(N))]$, which from Eqs. (138143) gives
$\tilde{W}^{\rho\sigma}_{n}[V\otimes Q]=\tilde{W}^{\rho}_{n}[V]+\tilde{W}^{\sigma}_{n}[Q].$ (145)
The one-dimensional stable symmetric winding number is an additive coset in
$\tilde{w}^{\rho}_{1}[V]\in\mathcal{R}^{-1}\mathrm{Rep}(G)[\pi_{1}(U(N))]/(1\cdot\pi_{1}(U(N))).$ (146)
Again, as $\pi_{n}(U(N))\in\{0,{\mathbb{Z}}\}$, these groups are always isomorphic to either the trivial group or $\mathcal{R}^{-1}R_{G}$.
Note that Eq. (142) implies that $\tilde{W}^{\rho_{1}}_{n}[V]=\tilde{W}^{\rho_{2}}_{n}[Q]$ if and only if there are representations $\sigma\otimes\rho_{1}$ and $\sigma\otimes\rho_{2}$ in $\mathcal{R}$ such that the unstable winding numbers
$W_{n}[V\otimes\mathbbm{1}_{\sigma\rho_{2}}]=\sigma\rho_{2}\cdot W_{n}[V]=\sigma\rho_{1}\cdot W_{n}[Q]=W_{n}[Q\otimes\mathbbm{1}_{\sigma\rho_{1}}]$ (147)
are equal. Thus, $V$ and $Q$ are strongly equivalent, exactly as we want.
We have identified a complete set of stable invariants among the elements of the localization $\mathcal{R}^{-1}R_{G}$, but not all elements in $\mathcal{R}^{-1}R_{G}$ necessarily correspond to achievable invariants with a given $\mathcal{R}$. The subset which can occur is that where the numerator involves irreps which also occur in the denominator. We define the additive subgroup of $\mathcal{R}^{-1}R_{G}$
$\mathcal{K}^{\mathcal{R}}=\Bigg{\{}\frac{\sum_{r\text{ irrep}}n_{r}r}{\rho}\in\mathcal{R}^{-1}R_{G}\,:\,n_{r}\neq 0\implies\\ \rho=s+r\text{ for some }s\in\mathrm{Rep}(G)\Bigg{\}}.$ (148)
This is precisely the subgroup of winding number invariants which can occur.
To conclude, the homotopy groups of symmetric zero-dimensional QCA are
$\pi_{n}(\mathrm{QCA}_{0}^{\mathcal{R}})\cong\left\{\begin{array}[]{l l}\leq X^{T}_{G}&\quad n=0,\\ \mathcal{K}^{\mathcal{R}}/{\mathbb{Z}}&\quad n=1,\\ 0&\quad n\geq 2\text{ even,}\\ \mathcal{K}^{\mathcal{R}}&\quad n\geq 3\text{ odd,}\end{array}\right.$ (149)
where by ${\mathbb{Z}}$ we mean the subgroup generated by $1\in\mathcal{K}^{\mathcal{R}}$. We have also not precisely characterized the restriction on $\pi_{0}(\mathrm{QCA}_{0}^{\mathcal{R}})\leq X_{G}^{T}$ imposed by the limited ancillae in $\mathcal{R}$, as our focus is on the less familiar higher homotopy groups.
As an example, we consider $G=U(1)$. Then all irreducible representations are one-dimensional, and given by $U(1)\ni e^{i\theta}\mapsto e^{im\theta}$ (in multiplicative notation). Identifying this irrep with the polynomial $x^{m}$, the representation ring is isomorphic to the ring of integer Laurent polynomials in $x$,
$R_{U(1)}\cong{\mathbb{Z}}[x,x^{-1}].$ (150)
We have that $X_{U(1)}\cong{\mathbb{Z}}$, so $X_{U(1)}^{T}\cong 0$ is trivial. Let us suppose that the monoid of ancillae $\mathcal{R}$ contains all finite dimensional representations of $U(1)$, so $\mathcal{R}$ is the full representation rig and thus isomorphic to Laurent polynomials with non-negative coefficients,
$\mathcal{R}\cong{\mathbb{N}}[x,x^{-1}].$ (151)
Then $\mathcal{K}^{\mathcal{R}}$ is the additive group of rational polynomials with positive denominators where all powers which occur in the numerator also occur in the denominator,
$\mathcal{K}^{\mathcal{R}}\cong\bigg{\{}\frac{p(x)}{q(x)}\in\mathrm{Frac}\,{\mathbb{Z}}[x]\,:\,q(x)\in{\mathbb{N}}[x],\\ x^{n}\in p(x)\implies x^{n}\in q(x)\bigg{\}}.$ (152)
We used that one can divide the numerator and denominator of a ratio of Laurent polynomials by the most negative power of $x$ in either to obtain a ratio of usual polynomials. Then the homotopy groups of $\mathrm{QCA}^{\mathcal{R}}_{0}$ with this set of ancillae is given by Eq. (149). We note that the rational polynomial structure which appeared in Ref. [26] is again occurring here, as it is a consequence of the representation theory of $U(1)$.
D.4 One dimension with symmetry
The study of the connected components $\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})$ has been considered in Refs. [93, 26, 114], where complete classifications were found for $U(1)$ and finite cyclic groups of prime order. In this appendix, we use the representation theory technology of the previous section to generalize these classifications to arbitrary topological groups $G$ and countable collections of ancillae representations $\mathcal{R}$.
D.4.1 Unstable invariants
Two one-dimensional QCA $V_{1}$ and $V_{2}$ are path equivalent if and only if $V_{1}V_{2}^{\dagger}$ is a finite depth circuit [58, 59]. Leveraging this fact simplifies the definition of invariants and proofs of classifications. Indeed, all one-dimensional QCA can be expressed as circuits of isometries [132]. That is, possibly after coarse graining by blocking sites together, any QCA $V$ on a one-dimensional chain with sites $x$ and local Hilbert space dimensions $N_{x}$ is equal to a product
$V=\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}^{\dagger}\right)\left(\prod_{x\in 2{\mathbb{Z}}}u_{x,x+1}\right)$ (153)
where
$u_{x,x+1}\in U(N_{x}N_{x+1}\to N^{\prime}_{x}N^{\prime}_{x+1}).$ (154)
That is, $V$ can be written as a finite depth circuit where the internal legs of the circuit do not necessarily have the same dimension as the external legs—the gates of the circuit are isometries, rather than unitaries. However, for the entire circuit to be unitary, we must have $N_{x}N_{x+1}=N^{\prime}_{x}N^{\prime}_{x+1}$.
Then $V$ commutes with the on-site symmetry $\prod_{x}\rho_{x}$ if and only if
$u_{x,x+1}\rho_{x}(g)\rho_{x+1}(g)=v_{x}(g)v_{x+1}(g)u_{x,x+1}$ (155)
for all $g\in G$ and $x\in{\mathbb{Z}}$, where $v_{x}(g)\in U(N^{\prime}_{x})$ is a unitary on the internal leg [132].
The product of two consecutive $v_{x}v_{x+1}$ is, by Eq. (155), unitarily equivalent to the representation $\rho_{x}\rho_{x+1}$. However, this only implies that $v_{x}$ and $v_{x+1}$ are projective representations of $G$. The classifications of Ref. [93, 26, 114] rely on the following observation: $V$ is equal to a finite depth circuit with symmetric unitary gates (which implies they are connected by a symmetric path) if and only if there are functions $\phi_{x}:G\to U(1)$ such that $e^{i\phi_{x}}v_{x}$ is unitarily equivalent to $\rho_{x}$ for all $x$. That is, the projective representation $v_{x}$ can be lifted to an ordinary unitary representation, and that representation is precisely $\rho_{x}$. A more concise statement of the condition is that $v_{x}$ and $\rho_{x}$ are isomorphic when both regarded as projective representations.
Indeed, suppose that
$e^{i\phi_{x}(g)}w_{x}v_{x}(g)w_{x}^{\dagger}=\rho_{x}(g)\quad\text{for all }g\in G.$ (156)
From the fact that $v_{x}v_{x+1}$ is unitarily equivalent to $\rho_{x}\rho_{x+1}$, we conclude that $\phi_{x}=-\phi_{x+1}$. Further, $N_{x}=N^{\prime}_{x}$ as $w_{x}$ is unitary, so we can write $V$ as a circuit of unitary gates
$V=\left(\prod_{x\in 2{\mathbb{Z}}+1}u^{\prime\dagger}_{x,x+1}\right)\left(\prod_{x\in 2{\mathbb{Z}}}u^{\prime}_{x,x+1}\right)$ (157)
with
$u^{\prime}_{x,x+1}=w_{x}w_{x+1}u_{x,x+1}\\ =(e^{i\phi_{x}}w_{x})(e^{i\phi_{x+1}}w_{x+1})u_{x,x+1}$ (158)
such that
$u^{\prime}_{x,x+1}\rho_{x}\rho_{x+1}$ $=w_{x}w_{x+1}v_{x}v_{x+1}u_{x,x+1}$ (159)
$=(e^{i\phi_{x}}w_{x}v_{x}w_{x}^{\dagger})(e^{i\phi_{x+1}}w_{x+1}v_{x+1}w_{x+1}^{\dagger})$
×$\qquad\qquad\qquad\times w_{x}w_{x+1}u_{x,x+1}$ (160)
$=\rho_{x}\rho_{x+1}u^{\prime}_{x,x+1}$ (161)
is symmetric.
Conversely, if $V$ can be written as a circuit with symmetric unitary gates, as in Eq. (157), then we must have that
$\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}u^{\prime\dagger}_{x,x+1}\right)\left(\prod_{x\in 2{\mathbb{Z}}}u^{\prime}_{x,x+1}u^{\dagger}_{x,x+1}\right)=\mathbbm{1}.$ (162)
We see by, for instance, the fact that the operator entanglement across the $(x,x+1)$ bond in this product must vanish, that
$u^{\prime}_{x,x+1}u^{\dagger}_{x,x+1}=w_{x}w_{x+1}$ (163)
must be a tensor product of isometries $w_{x}\in U(N_{x}\to N_{x}^{\prime})$. Further, as the whole circuit is the identity, we have $w_{x}^{\dagger}w_{x}=\mathbbm{1}$ and thus $N_{x}^{\prime}\geq N_{x}$ for all $x$. However, $N_{x}N_{x+1}=N^{\prime}_{x}N^{\prime}_{x+1}$, and these conditions can only be simultaneously satisfied if $N_{x}=N^{\prime}_{x}$ for all $x$, and thus $w_{x}$ is a unitary gate.
Then we have
$v_{x}v_{x+1}u_{x,x+1}$ $=u_{x,x+1}\rho_{x}\rho_{x+1}$ (164)
$=w_{x}w_{x+1}u^{\prime}_{x,x+1}\rho_{x}\rho_{x+1}$ (165)
$=w_{x}w_{x+1}\rho_{x}\rho_{x+1}u^{\prime}_{x,x+1}$ (166)
$=(w_{x}\rho_{x}w^{\dagger}_{x})(w_{x+1}\rho_{x+1}w^{\dagger}_{x+1})u_{x,x+1},$ (167)
so that $v_{x}=e^{-i\phi_{x}}w_{x}\rho_{x}w^{\dagger}_{x}$, as we needed.
Thus, an invariant for symmetric QCA in one dimension must somehow measure the difference between $v_{x}$ and $\rho_{x}$. We split this into two parts. First, we need $v_{x}$ to lift to a usual unitary representation. It is well known that the obstruction to lifting a projective unitary representation to a unitary representation is captured by group cohomology with coefficients in $U(1)$, so there is an
$\omega_{x}\in H^{2}(G;U(1))$ (168)
such that $v_{x}$ lifts to a unitary representation if and only if $\omega_{x}=0$. (We use an additive notation for $\omega_{x}$.) Further, $e^{-i\omega_{x}}v_{x}$ always lifts to a unitary representation, say $\tilde{v}_{x}$. The unitary representation $\tilde{v}_{x}$ is unique up to multiplication by a one dimensional representation $\chi_{x}:G\to U(1)$. Thus, we can identify the isomorphism class of $\tilde{v}_{x}$, with a coset in the representation rig,
$\tilde{v}_{x}X_{G}\in\mathrm{Rep}(G)/X_{G},$ (169)
where $X_{G}$ is the multiplicative subgroup of units (one dimensional representations) in $\mathrm{Rep}(G)$. Then $V$ is a symmetric circuit if and only if
$(\omega_{x},\tilde{v}_{x}X_{G})=(0,\rho_{x}X_{G})$ (170)
for all $x$.
D.4.2 Stable invariants
While the $\omega_{x}$ invariant is stable to stacking ancillae, the $\tilde{v}_{x}X_{G}$ invariant is not, and both apparently depend on position $x$. Stabilizing the $\tilde{v}_{x}X_{G}$ invariant through the algebraic localization construction fixes both of these issues.
We modify the $\tilde{v}_{x}X_{G}$ invariant to
$\frac{v_{x}}{\rho_{x}}X_{G}\in\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G},$ (171)
which we now regard as a multiplicative group via the multiplication in $\mathrm{Rep}(G)$.
The quotient by $X_{G}$ may refer to either the quotient by $X_{G}/1\subseteq\mathcal{R}^{-1}\mathrm{Rep}(G)$ or performing the $X_{G}$ quotient first, then the localization. The results are isomorphic.
When appending additional ancillae to site $x$ with representation $\sigma_{x}$, $V\otimes\mathbbm{1}$ acts trivially on the ancilla, so $v_{x}$ just becomes $v_{x}\otimes\sigma_{x}$. Then $v_{x}\sigma_{x}/\rho_{x}\sigma_{x}=v_{x}/\rho_{x}$ is unchanged.
Further, because $v_{x}v_{x+1}=\rho_{x}\rho_{x+1}\in\mathcal{R}$, we have
$\omega_{x}+\omega_{x+1}=0\quad\text{and}\quad\frac{\tilde{v}_{x}}{\rho_{x}}\frac{\tilde{v}_{x+1}}{\rho_{x+1}}=1.$ (172)
So the invariant for $x+1$ is just the inverse of that for $x$, and has no new information. Thus, we can just focus on $x=0$, and henceforth drop the $x$ index.
D.4.3 Classification
We have a complete set of invariants in $H^{2}(G;U(1))\times\mathcal{R}^{-1}\mathrm{Rep}(G)/X_{G}$. Again, not all $(\omega,\tilde{v}X_{G}/\rho)$ can actually be achieved with a fixed $\mathcal{R}$. We focus primarily on $\tilde{v}X_{G}/\rho$.
The only numerators $\tilde{v}$ which may occur in the invariant $\tilde{v}X_{G}/\rho$ are those which occur as tensor factors in $\rho$ for some $\rho\in\mathcal{R}$. That is, the numerators belong to the saturation
^=$\hat{\mathcal{R}}=\{v\in\mathrm{Rep}(G)\,:\,\exists\rho\in\mathrm{Rep}(G)\\ \text{such that }v\otimes\rho\in\mathcal{R}\}.$ (173)
Note that $\hat{\mathcal{R}}$ is closed under multiplication by $\mathcal{R}$, so the localization $\mathcal{R}^{-1}\hat{\mathcal{R}}$ is well defined and a multiplicative group under the natural multiplication operation of fractions.
All the elements of $\hat{\mathcal{R}}$ may occur as invariants. Suppose $v_{0}/\rho_{0}\in\mathcal{R}^{-1}\hat{\mathcal{R}}$. Then, as $v_{0}$ belongs to $\hat{\mathcal{R}}$, there are representations $v_{1}\in\mathrm{Rep}(G)$ and $\rho_{1}\in\mathcal{R}$ and a unitary $u_{0,1}^{\prime}$ such that
$v_{0}\otimes v_{1}u_{0,1}^{\prime}=u_{0,1}^{\prime}\rho_{1}.$ (174)
We tensor in $\rho_{0}$ to get
$v_{0}\otimes(v_{1}\otimes\rho_{0})u_{0,1}=u_{0,1}\rho_{0}\otimes\rho_{1}.$ (175)
Then we define a QCA $V$ which commutes with the representation $\prod_{x\in 2{\mathbb{Z}}}\rho_{0x}\otimes\rho_{1({x+1)}}$ by
$V=\left(\prod_{x\in 2{\mathbb{Z}}+1}u_{x,x+1}^{\dagger}\right)\left(\prod_{x\in 2{\mathbb{Z}}}u_{x,x+1}\right),$ (176)
where $u_{x,x+1}$ is a translation of $u_{0,1}$ from Eq. (175). The invariant for the QCA $V$ is, by construction, $(v_{0}X_{G})/\rho_{0}$, or its inverse $(v_{1}\rho_{0}X_{G})/\rho_{1}$.
With regard to the invariant described by Eq. (168), let us briefly address whether all elements of $H^{2}(G,U(1))$ can actually be achieved by a QCA [recall that this was not necessarily the case for the $H^{1}(G,U(1))$ invariant discussed in $d=0$, for infinite groups $G$]. This invariant characterizes the obstruction in lifting the projective representation $v_{x}$ to a unitary representation. For finite groups, all such invariants can be achieved. Taking the local Hilbert space to be the span of $|g\rangle$ with $g\in G$, and given any 2-coycle $\alpha:G\times G\to U(1)$, we define
$v_{0}(g)=\rho_{\mathrm{reg}}(g)D(g)$ (177)
where $\rho_{\mathrm{reg}}$ is the regular representation and $D(g)$ is the diagonal matrix with elements $\langle h|D(g)|h\rangle=\alpha(g,h)$. Then $v_{0}$ is known to be a projective representation with obstruction $[\alpha]\in H^{2}(G;U(1))$. By a similar construction to the previous paragraph, we can construct a QCA (indeed, a finite depth circuit) from the unitary gate $u_{0,1}$ which conjugate $\rho_{\mathrm{reg}}\otimes\rho_{\mathrm{reg}}$ to $v_{0}\otimes v_{1}$, where $v_{1}$ has obstruction $-[\alpha]$. This gate exists by the vanishing of the obstruction for $v_{0}\otimes v_{1}$. We leave open the precise restrictions on $\omega\in H^{2}(G;U(1))$ imposed by limited ancillae representations or by $G$ being infinite.
The full classifying group is then a subgroup
$\pi_{0}(\mathrm{QCA}_{1}^{\mathcal{R}})\leq H^{2}(G;U(1))\oplus\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G},$ (178)
with only the first summand possibly allowing for this to be a proper subgroup. That is, the projection to the second summand is surjective. We express the second summand with a logarithm just to make the group operation addition. The induced homomorphism from evaluation on a product state, $e_{0}^{*}$, just corresponds to projecting to the first summand, and thus reproduces the group cohomology classification of invertible states in one dimension [66, 93].
The case of $G=U(1)$ again serves as a simple example [26, 27]. We take $\mathcal{R}\cong{\mathbb{N}}[x,x^{-1}]$ to consist of all representations of $U(1)$. Then ^$\hat{\mathcal{R}}\cong{\mathbb{N}}[x,x^{-1}]$ as well, and the algebraic localization is isomorphic to the multiplicative group of rational polynomials with positive coefficients,
$\mathcal{R}^{-1}\hat{\mathcal{R}}\cong\mathrm{Frac}\,{\mathbb{N}}[x]^{\times}.$ (179)
The group of units $X_{U(1)}$ consists of monomials, $x^{m}$. Expressing the quotient in the additive notation, this is
$\log\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{U(1)}\cong\log\mathrm{Frac}\,{\mathbb{N}}[x]^{\times}/{\mathbb{Z}}\log x.$ (180)
As $H^{2}(U(1);U(1))\cong 0$ is trivial, this is the complete classification of $U(1)$-symmetric QCA in one dimension.
It has also been observed that including the regular representation for a finite group $G$ in $\mathcal{R}$, $\rho_{\mathrm{reg}}\in\mathcal{R}$, reduces the classification to just $H^{2}(G;U(1))\oplus\log{\mathbb{Q}}^{\times}$ [93]. Indeed, in the current presentation it is clear that this is because $R_{G}$ has zero divisors when $G$ is a finite group. If $\rho_{\mathrm{reg}}$ is the regular representation, then in the representation rig we have $\rho_{\mathrm{reg}}\sigma=\mathrm{dim}(\sigma)\rho_{\mathrm{reg}}$, and so
$\rho_{\mathrm{reg}}[\sigma\mathrm{dim}(\sigma^{\prime})-\sigma^{\prime}\mathrm{dim}(\sigma)]=0,$ (181)
for any $\sigma,\sigma^{\prime}\in\mathrm{Rep}(G)$. Thus, $\sigma/\sigma^{\prime}=\mathrm{dim}(\sigma)/\mathrm{dim}(\sigma^{\prime})\in\mathcal{R}^{-1}\hat{\mathcal{R}}$, and any two representations of the same dimension define the same element of the algebraic localization, which is then isomorphic to ratios of positive integers, $\mathcal{R}^{-1}\hat{\mathcal{R}}/X_{G}\cong{\mathbb{Q}}^{\times}$. (More correctly, this should itself be a localization of the natural numbers, $\mathrm{dim}(\mathcal{R})^{-1}\mathrm{dim}(\hat{\mathcal{R}})$.)
D.5 Functorality
Any classification of symmetric objects should, ideally, be compatible with forgetting the symmetry. That is, it should be possible to map the classification of symmetric objects into the classification of objects without symmetry in a consistent way. This is generalized by the notion of the classification being functorial [103].
Suppose that we have a homomorphism of topological groups
$\phi:H\to G,$ (182)
and a representation $\rho:G\to U(N)$ of $G$. Then $\rho\phi$ is a representation of $H$,
[…]
Further, we have that $(\rho_{1}\oplus\rho_{2})\phi\simeq\rho_{1}\phi\oplus\rho_{2}\phi$ and $(\rho_{1}\otimes\rho_{2})\phi\simeq\rho_{1}\phi\otimes\rho_{2}\phi$, so that $\rho\mapsto\rho\phi$ defines an induced rig homomorphism
$\phi^{*}_{\mathrm{Rep}}:\mathrm{Rep}(G)\to\mathrm{Rep}(H).$ (184)
Note that the order of $G$ and $H$ are reversed. The map is contravariant.
What’s more, any symmetric $V\in\mathrm{QCA}_{d}^{\mathcal{R}}$ (or more generally a QCA subject to some $G$ action, not necessarily symmetric) can also be regarded as being symmetric with respect to the $H$ action $\rho\phi$. Thus, we have a continuous inclusion map
$\phi^{*}_{\mathrm{QCA}}:\mathrm{QCA}_{d}^{\mathcal{R}}\to\mathrm{QCA}_{d}^{\mathcal{R}^{\prime}}$ (185)
where ϕRep()subscriptsuperscriptitalic-$\phi^{*}_{\mathrm{Rep}}(\mathcal{R})\subseteq\mathcal{R}^{\prime}$. In turn, this continuous map induces homomorphisms
$\phi^{*}_{n}:\pi_{n}(\mathrm{QCA}_{d}^{\mathcal{R}})\to\pi_{n}(\mathrm{QCA}_{d}^{\mathcal{R}^{\prime}}).$ (186)
In Appendices D.3 and D.4, we classified homotopy groups of symmetric QCA with $d-n\leq 1$. More concretely, we constructed isomorphisms between $\pi_{n}(\mathrm{QCA}^{\mathcal{R}}_{d})$ and certain groups defined in terms of $\mathrm{Rep}(G)$. One should want these isomorphisms to be functorial—the induced map $\phi^{*}_{n}$ should be reproduced by applying $\phi^{*}_{\mathrm{Rep}}$ (or more correctly the induced map on the localization or quotient) to the invariant in $\mathcal{R}^{-1}\mathrm{Rep}(G)$.
It is not difficult to see that this is the case. Indeed, as the map $\phi^{*}_{\mathrm{QCA}}$ does not change the QCA $V$, only the representation, carrying through the calculation of the invariant just involves replacing all the representations $\rho$ which occur with $\rho\phi$, which is precisely the action of $\phi^{*}_{\mathrm{Rep}}$.
An important example is completely forgetting the symmetry, corresponding to the inclusion of the identity, $\phi:1\to G$. Then we have that all representations reduce to the identity operator $\rho\phi=\mathbbm{1}_{\mathrm{dim}(\rho)}$. In the rig $\mathrm{Rep}(G)$, this corresponds to the dimension map,
$\mathrm{dim}:\rho\mapsto\mathrm{dim}(\rho),$ (187)
where the right hand side is to be interpreted as the sum of $\mathrm{dim}(\rho)\in{\mathbb{N}}$ copies of the trivial one-dimensional representation $1$. For $G=U(1)$, identifying $\mathrm{Rep}(U(1))$ with ${\mathbb{N}}[x,x^{-1}]$, the dimension map just corresponds to evaluating the rational polynomial invariant $f(x)$ at $x=1$ [26],
$\mathrm{dim}(f(x))=f(1).$ (188)
D.6 Lattice symmetries
It is believed that the generalized cohomology classification of invertible states also captures lattice symmetries and antiunitary symmetries [103, 104]. We also believe this to be true for QCA, and in this section we outline the structure that this should take. We leave detailed explorations to Ref. [107] and future work.
In the picture of Appendix C.3, a QCA on some $d$-dimensional orientable manifold $M$ can be identified by a map from $M$ to $\mathrm{QCA}_{d}$
$V\in\mathrm{QCA}_{M}\longleftrightarrow v:M\to\mathrm{QCA}_{d}.$ (189)
More precisely, we suspect $\mathrm{QCA}_{M}$ and $\left[M,\mathrm{QCA}_{d}\right]$ to be homotopy equivalent. Lattice symmetries can be treated in this picture as actions of the symmetry group $G$ on the base manifold $M$ [116]. Then, identifying the QCA $V$ with the map $v$, the QCA is symmetric if
$v(x)=\rho_{g}v(g\cdot x)\rho_{g}^{\dagger}$ (190)
for all $g\in G$ and $x\in M$. Then QCA with lattice symmetries would be fixed points of the mapping space
$\left[M,\mathrm{QCA}_{d}\right]^{\mathcal{R}}$ (191)
under the simultaneous action of $G$ on $\mathrm{QCA}_{d}$ and $M$. The action on $\mathrm{QCA}_{d}$ will generically not be on-site. For instance, a reflection should also act to reflect the QCA $v(x)$.
Abstractly, we are lead to equivariant Bredon cohomology, which classifies such fixed points of mapping spaces when the target space is an $\Omega$-spectrum [131]. The cohomology may, in general, need to be twisted when either $M$ is not orientable or, as would be more common for physics, the symmetry action is orientation reversing (involves reflections) or antiunitary.
Appendix E Models
The swindle map $S(v,t)$ is used in the main text and Appendix B as a theoretical tool. In fact, it also provides a useful construction for explicit models of $d$-dimensional driven (S)ALT phases, given a QCA (or family of QCA) $v$ in one dimension lower. If $v$ is nearest-neighbor, so is $S(v,t)$ (in the $\ell_{\infty}$ distance—it may involve diagonal couplings in the lattice). Even better, a slightly modified model only involves coupling two qubits at a time in one dimension. Thus, these models are well suited to experimental realization in a variety of quantum simulation platforms [63, 64, 65].
In Appendix E.1, we define the modified swindle map which only couples disjoint pairs of qubits at any time, and show how disorder may be introduced into the swindle map without spoiling solubility [24]. Then we explore the particular example of the quasiperiodic energy pump (QP pump) [31, 32, 53] in Appendix E.2. This is a one-dimensional $U(1)$-SALT phase with two drive tones which, thus far, has not been experimentally realized. The model we obtain through the swindle construction is particularly simple, paving the way for experiments on this phase.
E.1 Modified swindle
If $v\in\mathrm{QCA}_{d-1}$ has a range $r=1$ ($vAv^{\dagger}$ is localized within one site of $A$), then $S(v,t)$ also has range $1$. However, say in one dimension, it typically couples a site $j$ to both $j-1$ and $j+1$ simultaneously. This makes experimental implementation of the swindle models in some quantum simulation platforms more complicated. Further, the swindle model is translationally invariant, and so does not enjoy the robustness of disordered systems to perturbations.
We address both of these issues by defining a modified swindle map $\tilde{S}(v,t)$. We split the loop defining the swindle unitary into three segments, $t\in[0,\tfrac{1}{3})$, $t\in[\tfrac{1}{3},\tfrac{2}{3})$, and $t\in[\tfrac{2}{3},1]$. Within the first segment, sites coupled along even numbered bonds along the first dimension of the lattice are coupled, while odd numbered bonds are coupled in the second segment. The third segment introduces on-site disorder to the model.
We pick arbitrary smooth functions $\eta_{1,2,3}(t)$ which respectively map $t\in[0,\tfrac{1}{3}]$, $[\tfrac{1}{3},\tfrac{2}{3}]$, and $[\tfrac{2}{3},1]$ to $\eta\in[0,1]$. Sensible choices for these functions are monotonic increasing. We also choose disordered $(d-1)$-dimensional evolution operators $u_{j}^{\mathrm{dis}}(\eta)$, which are required to commute with any desired symmetry of the model. Then, on open boundary conditions with $L\in 2{\mathbb{Z}}$ sites along the first dimension, we have
$\tilde{S}(v,t)=\prod_{j=0}^{L/2-1}u_{2j,2j+1}(v,\eta_{1}(t))$ (192)
[where $u_{j,j+1}$ is defined in Eq. (7)] for $t\in[0,\tfrac{1}{3})$,
$\tilde{S}(v,t)=v_{0}\left(\prod_{j=1}^{L/2-1}u_{2j-1,2j}(v,\eta_{2}(t))v^{\dagger}_{2j-1}v_{2j}\right)v^{\dagger}_{L}$ (193)
for $t\in[\tfrac{1}{3},\tfrac{2}{3})$, and
$\tilde{S}(v,t)=u_{j}^{\mathrm{dis}}(\eta_{3}(t))v_{0}v^{\dagger}_{L}$ (194)
for $t\in[\tfrac{2}{3},1]$ (Fig. 3).
Away from the boundaries, $\tilde{S}$ acts as the identity at $t=\tfrac{2}{3}$. The LIOMs are then determined by $u_{j}^{\mathrm{dis}}$. At the edges, $\tilde{S}$ does not act as the identity, and instead acts as $v$ or $v^{\dagger}$.
The Hamiltonian associated to the unitary $\tilde{S}$ can be obtained by differentiation,
$H(t)=[i\partial_{t}\tilde{S}(v,t)]\tilde{S}(v,t)^{\dagger}.$ (195)
E.2 QP pump
Figure 5: (a) The power pumped into the $\theta$ drive in the state $\rho_{0-,(L-1)+}$ (211). The average value is quantized to $\omega W$, where $W$ is the winding number of the unitary used to construct the QP pump model and $\omega$ is the frequency of the $\theta$ drive. The power is only nonzero when $t\bmod 1\in[0,\tfrac{1}{3})$. (b) The quantized power appears as an average slope (red dashed) in the energy absorbed by the $\theta$ drive (blue). In fact, the modified swindle model has $E_{\pm}(t)=\mp\omega Wt$ at integer times. Parameters: $\omega=(1+\sqrt{5})/2$, $W=1$, $\eta_{1}=\sin^{2}(3\pi t/2)$.
We carry through the modified swindle construction explicitly for the QP pump. This phase is of particular interest, as its edge states can be used to prepare highly excited non-classical states in quantum cavities [133]. Thus, an experimental realization of this model is especially desirable.
Our model is based on a one-dimensional chain of qubits with open boundary conditions and length $L\in 2{\mathbb{Z}}$. The construction for odd numbers of qubits is essentially the same. The resulting model will be qualitatively similar to that in Refs. [134, 32].
The model has a $U(1)$ symmetry generated by
$Q=\sum_{j}Q_{j}=\sum_{j}\tfrac{1}{2}(\mathbbm{1}+\sigma^{z}_{j}),$ (196)
so we choose a path to the swap gate which preserves this symmetry. Indeed, we may take
$\mathrm{swap}_{j,j+1}(t)=\exp\left[-i\frac{2\pi}{8}t\left(\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}\right)\right]$ (197)
up to a global phase, where $\vec{\sigma}_{j}\cdot\vec{\sigma}_{j+1}=\sigma^{x}_{j}\sigma^{x}_{j+1}+\sigma^{y}_{j}\sigma^{y}_{j+1}+\sigma^{z}_{j}\sigma^{z}_{j+1}$.
The lower dimensional QCA to which the swindle map will be applied will actually be a loop of QCA in zero dimensions. That is, a loop of unitary matrices,
$v_{j}(\theta)=\exp\left[-i\frac{W}{2}\theta\sigma^{z}_{j}\right],$ (198)
where $W\in{\mathbb{Z}}$ is the winding number and $\theta\in[0,2\pi)$ parameterizes the loop. Note that all $W\in{\mathbb{Z}}$ give distinct winding numbers with $U(1)$ symmetry. Without this symmetry, all odd (even) winding numbers would be connected by homotopies (Appendix C.2).
As the model is quasiperiodically driven, we will take $\theta(t)=\omega t\bmod 2\pi$ to be a linear function of $t$ with $\omega\not\in{\mathbb{Q}}$ irrational.
We take the disorder in the model to be on-site and diagonal in the $\sigma^{z}$ basis,
$u^{\mathrm{dis}}_{j}(t)=\exp\left[-it\delta_{j}\sigma^{z}_{j}\right],$ (199)
where $\delta_{j}$ is, say, uniformly random in $[-\Delta,\Delta]$.
The final elements of the modified swindle model we need are the functions $\eta_{1,2,3}(t)$. These are essentially the pulse shapes which implement each unitary as a function of $t$, and their specific choice should be tailored to the experimental platform in mind. We leave them unspecified, beyond requiring that they are smooth, monotonic, and their derivatives vanish at their end points. This ensures that there are no discontinuities in the pulses, which also assists in stabilizing localization [28, 99].
The Hamiltonian is then defined piecewise as
$H(t,\theta)=\left\{\begin{array}[]{l l}\omega h^{1}(\eta_{1}(t),\theta)+\dot{\eta}_{1}(t)h^{2}(\eta_{1}(t),\theta),&t\in[0,\tfrac{1}{3}),\\ \omega h^{3}(\eta_{2}(t),\theta)+\dot{\eta}_{2}(t)h^{4}(\eta_{2}(t),\theta),&t\in[\tfrac{1}{3},\tfrac{2}{3}),\\ \omega h^{5}+\dot{\eta}_{3}(t)h^{6},&t\in[\tfrac{2}{3},1],\end{array}\right.$ (200)
where $\dot{\eta}_{k}=\partial_{t}\eta_{k}$ and $H(t,\theta)$ has period $T=1$ in $t$ and period $2\pi$ in $\theta$. Defining $\sigma^{\pm}_{j}=\sigma_{j}^{x}\pm i\sigma_{j}^{y}$, the individual terms for the first pulse are
$h^{1}(\eta_{1},\theta)$ $=\sum_{j=0}^{L/2-1}h^{1}_{2j,2j+1}(\eta_{1},\theta)$ (201)
$h^{2}(\eta_{1},\theta)$ $=\sum_{j=0}^{L/2-1}h^{2}_{2j,2j+1}(\eta_{1},\theta),$ (202)
where
[…]
For the second pulse, we have
$h^{3}(\eta_{2},\theta)=\frac{W}{2}(\sigma^{z}_{0}-\sigma^{z}_{L-1})+\sum_{j=1}^{L/2-1}h^{3}_{2j-1,2j}(\eta_{2},\theta),\quad\text{and}\quad h^{4}(\eta_{2},\theta)=\sum_{j=1}^{L/2-1}h^{2}_{2j-1,2j}(\eta_{2},\theta),$ (205)
where
$h_{j,j+1}^{3}(\eta_{2},\theta)=\frac{W}{4}\bigg{[}[1+\cos(W\theta)+\cos(2\pi\eta_{2})(1-\cos(W\theta))](\sigma^{z}_{j}-\sigma^{z}_{j+1})\\ -\frac{i}{2}e^{-iW\theta}[\sin(\pi\eta_{2})(1-2i\sin(W\theta))+\sin(2\pi\eta_{2})(1-\cos(W\theta))]\sigma_{j}^{+}\sigma_{j+1}^{-}+\text{H.c.}\bigg{]}.$ (206)
Finally, the disorder pulse has
$h^{5}=\frac{W}{2}(\sigma^{z}_{0}-\sigma^{z}_{L-1}),\quad\text{and}\quad h^{6}=\sum_{j=0}^{L-1}\delta_{j}\sigma^{z}_{j}.$ (207)
Additionally, rotating frame transformations generated by combinations of $\sigma^{z}_{j}$ can be used to eliminate some of the complex time dependence of the coupling terms. For instance, making the transformation
$\tilde{S}(t)\mapsto e^{i\omega t\frac{W}{2}\sum_{j}j\sigma^{z}_{j}}\tilde{S}(t)$ (208)
eliminates the $e^{-iW\theta}$ factors in all the expressions above, at the cost of introducing a time-independent term
$H^{\prime}=-\omega\frac{W}{2}\sum_{j=0}^{L-1}j\sigma^{z}_{j}$ (209)
to the Hamiltonian. That is, the time-dependent vector potential in the model can be traded for a linear scalar potential. On the other hand, if implementing time-dependent detunings $\sigma^{z}_{j}-\sigma^{z}_{j+1}$ is more challenging in some architecture, moving to a frame co-rotating with those terms of the Hamiltonian moves this time dependence to the $\sigma^{+}_{j}\sigma^{-}_{j+1}$ exchange term.
The main interesting feature of the QP pump is its edge states, which transfer energy from one drive to the other. To compute the pumped power in the edge state, we evaluate the expectation value of
$P(t,\theta)=-\omega\partial_{\theta}H(t,\theta),$ (210)
which is the work done by the qubit chain on the drive with phase $\theta$. We evaluate this operator in a mixture of quasienergy states, with the initial density matrix being
$\rho_{0\pm,(L-1)\pm}=\frac{1}{2^{L}}(\mathbbm{1}\pm\sigma^{z}_{0})(\mathbbm{1}\pm\sigma^{z}_{L-1}).$ (211)
This is maximally mixed in the bulk LIOM states, but has a definite state at both edges. The expectation value of the power operator is thus
$\langle P(t,\theta)\rangle_{\pm\pm}=\mathrm{Tr}\left[P\tilde{S}\rho_{0\pm,(L-1)\pm}\tilde{S}^{\dagger}\right]$ (212)
where we suppressed the dependence on $t$ and $\theta$ in the right hand side.
The model is constructed so that $\tau_{j}(t=0)=\sigma^{z}_{j}$, and the Heisenberg picture operator $\sigma^{z\mathrm{H}}_{j}(t)$ is a function of ${\vec{\theta}}$. We can thus find the edge LIOMs within a period by “reverse Heisenberg evolution” [28] of the $\sigma^{z}_{0}$ operator. We have, for $t\in[0,\tfrac{1}{3})$
$\tau^{z}_{0}(t,\theta)=u_{0,1}(v(\theta),\eta_{1}(t))\sigma^{z}_{0}u_{0,1}^{\dagger}(v(\theta),\eta_{1}(t))$ (213)
and $\tau^{z}_{0}=\sigma^{z}_{0}$ otherwise. Equation (212) then gives that
$\langle P(t,\theta)\rangle_{\pm\pm}=\frac{1}{2^{L}}\mathrm{Tr}\left[P(\mathbbm{1}\pm\tau^{z}_{0})(\mathbbm{1}\pm\tau^{z}_{L-1})\right].$ (214)
Further using that the power is a traceless local operator, $P=\sum_{j}P_{j,j+1}$, and that $\tau^{z}_{0,L-1}$ have disjoint support, we can expand Eq. (214) as
$\langle P(t,\theta)\rangle_{\pm\pm}=\pm\frac{1}{2^{L}}\mathrm{Tr}\left[P\tau^{z}_{0}\right]\pm\frac{1}{2^{L}}\mathrm{Tr}\left[P\tau^{z}_{L-1}\right].$ (215)
Calculating the LIOMs from the definition of $u_{0,1}$ gives
$u_{0,1}(v(\theta),\eta_{1})\sigma^{z}_{0}u_{0,1}^{\dagger}(v(\theta),\eta_{1})=\frac{1}{2}\bigg{[}\sigma^{z}_{0}+\sigma^{z}_{1}+\frac{1}{2}[1+\cos(W\theta)+\cos(2\pi\eta_{1})(1-\cos(W\theta))](\sigma^{z}_{0}-\sigma^{z}_{1})\\ +\frac{e^{-iW\theta}}{4}[2\sin(\pi\eta_{1})\sin(W\theta_{1})+i\sin(2\pi\eta_{1})(1-\cos(W\theta))]\sigma^{+}_{0}\sigma^{-}_{1}+\text{H.c.}\bigg{]}.$ (216)
Similarly, $\tau^{z}_{L-1}(t,\theta)$ is given by $\sigma^{z}_{L-1}$ except during the first pulse, which is determined by
$u_{L-2,L-1}(v(\theta),\eta_{1})\sigma^{z}_{L-1}u_{L-2,L-1}^{\dagger}(v(\theta),\eta_{1})=\frac{1}{2}\bigg{[}\sigma^{z}_{L-2}+\sigma^{z}_{L-1}-\frac{1}{2}[1+\cos(W\theta)+\cos(2\pi\eta_{1})(1-\cos(W\theta))](\sigma^{z}_{L-2}-\sigma^{z}_{L-1})\\ -\frac{e^{-iW\theta}}{4}[2\sin(\pi\eta_{1})\sin(W\theta_{1})+i\sin(2\pi\eta_{1})(1-\cos(W\theta))]\sigma^{+}_{0}\sigma^{-}_{1}+\text{H.c.}\bigg{]}.$ (217)
The power pumped by the left (smaller $j$) edge state is
$\frac{1}{2^{L}}\mathrm{Tr}\left[P\tau^{z}_{0}\right]=-\frac{\omega}{2^{L}}\mathrm{Tr}\left[(\omega\partial_{\theta}h^{1}_{0,1}+\dot{\eta}_{1}\partial_{\theta}h^{2}_{0,1})\tau^{z}_{0}\right]=\frac{\omega W}{4}\bigg{[}\frac{\omega W}{2}\sin(W\theta)(1-\cos(2\pi\eta_{1}))\\ +\pi\dot{\eta}_{1}(\sin(2\pi\eta_{1})(1-\cos(W\theta))-\sin(\pi\eta_{1}))\bigg{]},$ (218)
for $t\in[0,\tfrac{1}{3})$, and $\mathrm{Tr}\left[P\tau^{z}_{0}\right]=0$ otherwise. Further, $\mathrm{Tr}\left[P\tau^{z}_{0}\right]=-\mathrm{Tr}\left[P\tau^{z}_{L-1}\right]$, so the pumped power is nonzero only when the two edge states are opposite. We will subsequently let $\pm$ refer only to the occupation of the LIOM at site $j=0$, and assume the other edge LIOM is in the opposite state. Then the total pumped power is
$\langle P(t,\theta)\rangle_{\pm}=\pm\frac{\omega W}{2}\bigg{[}\frac{\omega W}{2}\sin(W\theta)\left(1-\cos(2\pi\eta_{1}(t))\right)+\pi\dot{\eta}_{1}(t)\left(\sin(2\pi\eta_{1}(t))(1-\cos(W\theta))-\sin(\pi\eta_{1}(t))\right)\bigg{]}$ (219)
for $t\in[0,\tfrac{1}{3})$. This function is plotted as a function of $t$ in Fig. 5.
We can compute the average pumped power by integrating $\langle P(t,\theta)\rangle_{\pm}$ over $t$ and $\theta$. This gives
$\bar{P}_{\pm}=\frac{1}{2\pi}\int_{0}^{2\pi}\mathrm{d}\theta\int_{0}^{1}\mathrm{d}t\,\langle P(t,\theta)\rangle_{\pm}=\mp\omega W.$ (220)
That is, $W$ photons of frequency $\omega$ are pumped into (or out of, if $W<0$) the $\theta$ drive per period of $t$ (which we have set to 1) when $\langle\tau^{z}_{0}\rangle=-1$ and $\langle\tau^{z}_{L-1}\rangle=+1$.
E.2.1 Invariant
Finally, we briefly explain where this instance of the QP pump sits in the classification of Appendix D.3. Specifically, we find its associated invariant.
The QP pump is $U(1)$-symmetric and has $d=1$, $n=2$. It thus has the same classifying group as loops of unitaries in $d=0$ with $U(1)$ symmetry. Indeed, this is the object from which our model was constructed. From Appendix D.3, to get the classification of $v(\theta)$ we must find its one-dimensional winding number in each symmetry sector. We have
$v(\theta)=e^{-iW\theta/2}\begin{pmatrix}1&0\\ 0&e^{iW\theta}\end{pmatrix},\quad Q_{0}=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}.$ (221)
Thus, up to a global phase, $v(\theta)$ has winding number $W$ in the charge 0 sector, and winding number 00 in the charge 1 sector.
The invariant associated to $v(\theta)$ is a coset [Eq. (141)] of formal ratios of representations with the numerator weighted by the winding numbers. We identify the charge 0 representation with 1 and the charge 1 representation with $x\in{\mathbb{N}}[x,x^{-1}]\cong\mathrm{Rep}(U(1))$. Then the invariant is
$\tilde{w}_{1}^{Q}[v]=\frac{1\cdot W+x\cdot 0}{1+x}+{\mathbb{Z}}=\frac{W}{1+x}+{\mathbb{Z}}.$ (222)
We see that all distinct values of $W$ indeed define inequivalent values of the invariant.
Suppose now that we explicitly broke the $U(1)$ symmetry. Then the charge sector $1\leftrightarrow x$ should be mapped to having charge 00. The new invariant can be found by evaluating the rational function $\tilde{w}_{1}^{Q}[v]$ at $x=1$ (Appendix D.5). We find
$\tilde{w}_{1}^{Q}[v]\mapsto\frac{W}{2}+{\mathbb{Z}}\in{\mathbb{Q}}/{\mathbb{Z}}.$ (223)
Now, all even $W$ are equivalent to the trivial phase. Interestingly, the odd $W$ phases remain non-trivial. They no longer support a quantized edge current, as we can deform the model with $W=+1$ to $W=-1$, but there is apparently still some nontrivial dynamics at the edge.
Indeed, we believe the odd $W$ phase without symmetry to be the two-tone-driven phase of Refs. [29, 30]. While the solvable limit of the model in Ref. [29] is highly singular, which makes computing invariants from integral formulae inconvenient, the edge phenomenology described in Ref. [29] is consistent with what we expect from having $\tilde{w}_{1}\neq 0\bmod 1$. Namely, the edge dynamics was argued to be anomalous in that it could not be achieved by a zero-dimensional localized [in the sense that a continuous micromotion operator exists as in Eq. (35)] system which did not occur at the edge of an extended one-dimensional system. As the model in Ref. [29] is based around a version of the AKLT model, which can be trivialized when stacked with itself, it is thus reasonable to suppose that $\tilde{w}_{1}=1/2\bmod 1$ for that model, as this is the only element of ${\mathbb{Q}}/{\mathbb{Z}}$ of order 2. Thus, the abstract classification reveals a relationship between the interacting QP pump and Refs. [29, 30] which was not clear previously. It also extends the arguments for stability in Ref. [29] to the nonperturbative regime, where a picture of emergent symmetry (as employed in that reference) becomes inapplicable.